### v1

# Sonnet: Spectral Operator Neural Network for Multivariable Time Series Forecasting

Abstract
Multivariable time series forecasting methods can integrate information from exogenous variables, leading to significant prediction accuracy gains. Transformer architecture has been widely applied in various time series forecasting models due to its ability to capture long-range sequential dependencies. However, a naïve application of transformers often struggles to effectively model complex relationships among variables over time. To mitigate against this, we propose a novel architecture, namely the Spectral Operator Neural Network (Sonnet). Sonnet applies learnable wavelet transformations to the input and incorporates spectral analysis using the Koopman operator. Its predictive skill relies on the Multivariable Coherence Attention (MVCA), an operation that leverages spectral coherence to model variable dependencies. Our empirical analysis shows that Sonnet yields the best performance on $34$ out of $47$ forecasting tasks with an average mean absolute error (MAE) reduction of $1.1\%$ against the most competitive baseline (different per task). We further show that MVCA—when put in place of the naïve attention used in various deep learning models—can remedy its deficiencies, reducing MAE by $10.7\%$ on average in the most challenging forecasting tasks.
Source code: github.com/ClaudiaShu/Sonnet
1 Introduction
Multivariable time series (MTS) forecasting methods learn from multiple input variables to predict a single target variable [17]. MTS models are deployed in many real-life applications, such as financial modelling [35, 3, 46, 54], numerical weather prediction [13, 55, 8], and computational epidemiology [11, 42, 7, 31]. Recent machine learning models for time series forecasting tend to emphasise multivariate formulations [34, 58, 24, 2, 9]. While such models provide important insights for integrating traditional time series forecasting into deep learning techniques—e.g. capturing seasonality [23], frequency-domain methods [62, 63], and autoregression [58, 34]—they often overlook the benefits of leveraging external information through exogenous indicators.
In fact, empirical evidence suggests that capturing dependencies from external variables can increase the risk of overfitting for complex models [34]. Moreover, existing methods that capture inter-variable dependencies [60, 51] have, in certain forecasting tasks, been outperformed by models that do not [23, 26]. However, this can be attributed to the choice of forecasting benchmarks, which, in many occasions, are characterised by strong seasonal patterns [23] (see also Appendix E.4), that can be effectively modelled using simpler autoregressive methods [58, 23]. Additionally, models that do not capture inter-variable dependencies benefit from learning using longer look-back windows without significantly increasing modelling complexity and computational cost [15], making them less likely to overfit. Nevertheless, these empirical outcomes are not necessarily valid for multivariable forecasting scenarios, and especially for tasks where exogenous variables have strong predictive power. In such cases, explicitly modelling interactions across variables facilitates better performance [52, 43].
Prior work has explored various ways in capturing inter-variable dependencies for multivariate or MTS forecasting tasks. Methods applying attention [48] across variables [25, 18] can capture nonlinear dependencies, but disrupt temporal information by embedding sequences along the time dimension. Crossformer [60] attempts to rectify this by proposing a modified transformer structure where input is split into time series patches allowing to capture dependencies across both the time and variable dimensions at the sub-sequence level. ModernTCN [27], on the other hand, uses a convolutional kernel over both the variable and time dimension to jointly capture the inter- and intra-variable dependencies. However, both Crossformer and ModernTCN suffer from GPU computational overhead as the number of exogenous variables or the length of the input time series increases [43, 64]. TimeXer [52] models dependencies between exogenous variables and the target variable separately, and then joins their learned embeddings with a cross-attention module. DeformTime [43] yields superior results by using deformable attention to incorporate information from exogenous variables at different time steps. However, it only captures inter-variable dependencies within the reception field of a convolutional kernel, which limits the exploration of a wider range of exogenous variables.
Frequency-based analysis, including using Fourier [6, 44] or wavelet transform [47, 59, 4, 12, 56], has been widely used in statistical approaches for identifying periodic patterns [37] as well as in machine learning methods for forecasting [20, 62, 63, 36]. These methods compress temporal information (due to the Fourier transform) while some of them focus on capturing intra-variable dependencies [62, 63]. Wavelet transform, albeit dependent on the chosen mother wavelet [10, 32], can maintain both time and frequency information. In the frequency domain, spectral coherence [53] serves as a powerful tool for capturing the correlation between variables at different frequencies [45, 30]. To improve input projections into the frequency domain, Lange et al. [20] used a Koopman operator [29, 40, 5, 24], i.e. a spectral function that enables linear modelling of nonlinear changes to better address nonlinearities. More recently, AdaWaveNet [56] decomposed the input series into seasonal and long-term components, and used the wavelet transform to capture periodic information. This method accounts for inter-variable dependencies only over the seasonal components.
Motivated by the aforementioned remarks, we propose Spectral Operator Neural Network (Sonnet), a model that captures time series dependencies in the spectral domain using a learnable wavelet transform. Sonnet captures both intra- and inter-variable dependencies using a novel frequency-domain Multivariable Coherence Attention (MVCA) layer. Furthermore, it deploys a learnable Koopman operator that facilitates linearised transitions of temporal states [22]. MVCA demonstrates stand-alone effectiveness when integrated into existing architectures, outperforming naïve attention and other modified attention mechanisms. Our contributions can be summarised as follows:
  1. 1.
    We propose Sonnet, a novel neural network architecture for MTS forecasting that captures inter-variable dependencies via adaptable time-frequency spectral operators while enforcing stability through learnable Koopman dynamics.
  2. 2.
    We introduce MVCA, an attention mechanism designed to model interactions between variables by leveraging their spectral coherence, a frequency-domain measurement of dependency. Unlike conventional self-attention, which computes pairwise similarity via dot products, MVCA captures temporal relationships through their cross-spectral density with the inclusion of frequency information from all variables, enhancing variable dependency modelling for MTS tasks.
  3. 3.
    We assess forecasting accuracy on carefully curated MTS data sets. This includes established benchmarks complemented by tasks on weather forecasting, influenza prevalence, electricity consumption, and energy prices. The weather prediction and influenza prevalence tasks support a more thorough evaluation because they include a substantial amount of exogenous variables, multiple years ($\geq 10$) for training, and multiple test periods ($\geq 3$).
  4. 4.
    Sonnet reduces mean absolute error (MAE) by $1.1\%$ on average compared to the most performant baseline (varies each time). In the more challenging tasks of influenza and weather modelling, MAE is reduced by $3.5\%$ and $2\%$, respectively. Performance gains persist as the forecasting horizon increases, demonstrating the effectiveness of Sonnet in longer-term forecasting.
2 MTS forecasting task definition
We focus exclusively on multivariable time series (MTS) forecasting, whereby multiple input variables are used to predict a single target variable. We note that although for some baseline models multiple output variables may be present (multivariate forecasting), our evaluation is restricted to the prediction of one specific output. All models are trained and evaluated under a rolling window setup, with a look-back window and a forecasting horizon of $L$ and $H$ time steps, respectively. Specifically, at each time step $t$, the $C$ observed exogenous variables over $L$ past time steps, $\{t-L+1,\dots,t-1,t\}$, are captured in an input matrix $\mathbf{X}_{t}\in\mathbb{R}^{L\times C}$. The autoregressive signal for the target (endogenous) variable is denoted by $\mathbf{y}_{t-\delta}\in\mathbb{R}^{L}$; this encompasses time steps $\{t-\delta-L+1,\dots,t-\delta-1,t-\delta\}$, where $\delta\in\mathbb{N}_{0}$ is an optional delay applied when the target variable is observed with a temporal lag. We capture the exogenous and endogenous input variables in $\mathbf{Z}_{t}=[\mathbf{X}_{t},\mathbf{y}_{t-\delta}]\in\mathbb{R}^{L\times(C+1)}$. The goal is to predict the target variable at time step $t+H$, where $H$ denotes the forecasting horizon. Hence, the output of a forecasting model is denoted by $\mathbf{y}_{t+H}\in\mathbb{R}^{H}$ and holds forecasts for time steps $\{t+1,\dots,t+H-1,t+H\}$. For models that conduct multivariate forecasting, the output includes predictions for all covariates and hence is denoted by $\mathbf{Y}_{t+H}=[\mathbf{X}_{t+H},\mathbf{y}_{t+H}]\in\mathbb{R}^{H\times(C+1)}$. The forecasting task is to learn $f:\mathbf{Z}_{t}\rightarrow\mathbf{y}_{t+H}$ or $\mathbf{Y}_{t+H}$. Performance is measured based on the endogenous forecast at time step $t+H$, i.e. the last (temporally) element of $\mathbf{y}_{t+H}$, $y_{t+H}\in\mathbb{R}$. For notational simplicity, we omit temporal subscripts and use $\mathbf{Z}$ for $\mathbf{Z}_{t}$, $\mathbf{y}$ for $\mathbf{y}_{t-\delta}$, and $\mathbf{X}$ for $\mathbf{X}_{t}$.
3 Spectral coherence with Sonnet
In this section, we provide a detailed description of our proposed model, Sonnet. It operates in the spectral domain via a learnable wavelet transform and uses a Multivariable Coherence Attention (MVCA) module to capture both inter- and intra-variable dependencies. We further use a Koopman projection layer that enables stable temporal evolution via a learned linear operator.
3.1 Joint embedding of exogenous and endogenous variables
Given an input matrix $\mathbf{Z}\in\mathbb{R}^{L\times(C+1)}$ that consists of the exogenous variables $\mathbf{X}\in\mathbb{R}^{L\times C}$ and endogenous variable $\mathbf{y}\in\mathbb{R}^{L}$, we first obtain the embeddings of $\mathbf{X}$ and $\mathbf{y}$ independently. These are denoted by $\mathbf{E}_{x}\in\mathbb{R}^{L\times\alpha d}$ and $\mathbf{E}_{y}\in\mathbb{R}^{L\times\left(1-\alpha\right)d}$. They are derived using learnable weight matrices $\mathbf{W}_{x}\in\mathbb{R}^{C\times\alpha d}$ and $\mathbf{W}_{y}\in\mathbb{R}^{1\times\left(1-\alpha\right)d}$, as in $\mathbf{E}_{x}=\mathbf{X}\mathbf{W}_{x}$, where $d$ is the embedding dimension and $\alpha\in[0,1]$ a hyperparameter that controls the projected dimensionality of $\mathbf{X}$ and $\mathbf{y}$ in the final embedding. [^1] We make sure that both products $\alpha d$ and $(1-\alpha)d\in\mathbb{N}$ to avoid dimension mismatch caused by rounding. For $\alpha=0$, the forecasting is based on the historical values of the target variable only, collapsing to an autoregressive setting. For $\alpha=1$, forecasting depends entirely on the exogenous variables. By concatenating along the feature dimension, we obtain the final embedding $\mathbf{E}=[\mathbf{E}_{x},\mathbf{E}_{y}]\in\mathbb{R}^{L\times d}$.
3.2 Learnable wavelet transform
We then transform the time series embedding into the wavelet space, i.e. a space that contains both time and frequency information. Specifically, after obtaining the input time series embedding $\mathbf{E}$, we first define a set of $K$ learnable wavelet transformations (also referred to as atoms), with the $k$-th atom held in a matrix $\mathbf{M}_{k}\in\mathbb{R}^{d\times L}$ derived by
$\mathbf{M}_{k}=\exp\left(-\mathbf{w}_{\alpha}\mathbf{t}^{2}\right)\times\cos\left(\mathbf{w}_{\beta}\mathbf{t}+\mathbf{w}_{\gamma}\mathbf{t}^{2}\right)\,\,,$ (1)
where $\mathbf{t}\in\mathbb{R}^{L}$ is a row vector capturing normalised time steps with $\mathbf{t}_{i}=i/(L-1)$ for $i=0,\dots,L-1$, and $\mathbf{w}_{\alpha}$, $\mathbf{w}_{\beta}$, and $\mathbf{w}_{\gamma}\in\mathbb{R}^{d}$ are learnable weight vectors that control the shape of the wavelet, each initialised randomly from a normal distribution. In particular, $\mathbf{w}_{\alpha}$ controls the width of the Gaussian envelope, and $\mathbf{w}_{\beta}$, $\mathbf{w}_{\gamma}$ respectively determine the linear or quadratic frequency modulation of the generated cosine waveforms. This formulation enables the atoms to adapt to localised time-frequency structures in the data. The time series embedding $\mathbf{E}\in\mathbb{R}^{L\times d}$ is then transformed into the wavelet space by projecting it onto each of the wavelet atoms $\mathbf{M}_{k}\in\mathbb{R}^{d\times L}$ using $\mathbf{P}_{k}=\mathbf{E}\odot\mathbf{M}_{k}^{\top}$, where $\mathbf{P}_{k}\in\mathbb{R}^{L\times d}$ denotes the embedding’s projection for the $k$-th atom. The transformed wavelet across all $K$ atoms is denoted by $\mathbf{P}\in\mathbb{R}^{K\times L\times d}$. The aforementioned steps help to preserve temporal structure while decomposing the input into multi-resolution time-frequency components using adaptive wavelet transforms that can capture both short- and long-term patterns in the data.
3.3 Multivariable Coherence Attention (MVCA)
In MTS forecasting, input variables can be both auto-correlated (to their own past values) and cross-correlated (to each other). To improve learning from these correlations, we propose MVCA, a module that supplements the standard attention. MVCA can capture inter- and intra-variable dependencies within the frequency domain using spectral density coherence. Its premise is that variables with a higher spectral coherence should contribute more to the attention output. MVCA can be used in place of any naïve attention module used in a forecasting method.
Given the obtained input embedding matrix in wavelet space $\mathbf{P}\in\mathbb{R}^{K\times L\times d}$, where $L$ and $d$ are the time and variable dimension, $K$ is the number of wavelets after transformation, we first linearly project it to query, key and value embeddings, denoted as $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}\in\mathbb{R}^{K\times L\times d}$, using weight matrices $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}\in\mathbb{R}^{d\times d}$ respectively, as in $\mathbf{Q}=\mathbf{P}\mathbf{W}_{q}$. We then consider the embeddings from different wavelet transformations (indexed by $k$) as separate attention heads. Specifically, for each attention head, we obtain the sub-tensor of $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$, denoted as $\mathbf{Q}_{h}$, $\mathbf{K}_{h}$, and $\mathbf{V}_{h}\in\mathbb{R}^{L\times d}$, as the query, key, and value embeddings. Therefore, each head captures a distinct subspace of the original embedding, where subspaces correspond to different wavelet transformations of the input. Including a multi-head structure enables the model to learn diverse dependencies in parallel [48].
We then apply a Fast Fourier Transform (FFT) along the variable dimension of each transformer head to transfer the query and key embeddings to the frequency domain, i.e. $\mathbf{Q}_{f}=\texttt{FFT}\left(\mathbf{Q}_{h}\right)$ and $\mathbf{K}_{f}=\texttt{FFT}\left(\mathbf{K}_{h}\right)$, with both $\mathbf{Q}_{f}$ and $\mathbf{K}_{f}\in\mathbb{C}^{L\times\ell}$, where $\ell=\lfloor\frac{d}{2}\rfloor+1$. Each frequency bin in the transformed data ($\mathbf{Q}_{f}$, $\mathbf{K}_{f}$) contains information from all original inputs to the FFT [21], capturing both fine-grained (high-frequency) and global (low-frequency) patterns. The cross-spectral density $\mathbf{P}_{qk}\,\in\,\mathbb{C}^{L\times\ell}$ and power-spectral densities $\mathbf{P}_{qq}$, $\mathbf{P}_{kk}\,\in\,\mathbb{R}^{L\times\ell}$ are obtained as follows:
$\mathbf{P}_{qk}=\mathbf{Q}_{f}\odot\mathbf{K}_{f}^{\ast}\,\,,\quad\mathbf{P}_{qq}=\mathbf{Q}_{f}\odot\mathbf{Q}_{f}^{\ast}\,\,,\quad\text{and}\,\quad\mathbf{P}_{kk}=\mathbf{K}_{f}\odot\mathbf{K}_{f}^{\ast}\,\,,$ (2)
where ‘’ denotes the complex conjugate. We average along the second dimension to obtain $\overline{\mathbf{P}}_{qk}\in\mathbb{C}^{L}$, $\overline{\mathbf{P}}_{qq}$, and $\overline{\mathbf{P}}_{kk}\in\mathbb{R}^{L}$ (see also Appendix B). The normalised spectral coherence $\mathbf{C}_{qk}\in\mathbb{R}^{L}$ is then computed using
$\mathbf{C}_{qk}=|\overline{\mathbf{P}}_{qk}|^{2}/\left(\overline{\mathbf{P}}_{qq}\cdot\overline{\mathbf{P}}_{kk}+\epsilon\right),$ (3)
where $\epsilon=10^{-6}$ mitigates division by 00. $\mathbf{C}_{qk}$ captures the linear dependency between sequences across frequency bands. Therefore, the coherence here assigns higher importance to the time steps where the query and key hold more similar averaged values across multiple frequencies.
Following the common design of attention layers [48], we scale, normalise, and regularise $\mathbf{C}_{qk}$, i.e. $\mathbf{A}_{h}=\texttt{Dropout}(\texttt{Softmax}(\mathbf{C}_{qk}/\sqrt{d}))$. The attention weights $\mathbf{A}_{h}\in\mathbb{R}^{L}$ of each head are first broadcast along the feature dimension to form $\mathbf{A}\in\mathbb{R}^{L\times d}$, which is then multiplied with the value representations $\mathbf{V}_{h}$ (element-wise) to produce head-specific outputs, $\mathbf{O}_{h}\in\mathbb{R}^{L\times d}=\mathbf{A}\odot\mathbf{V}_{h}$. We concatenate these outputs across all heads to obtain $\mathbf{O}_{r}\in\mathbb{R}^{K\times L\times d}$. We then use a $2$-layer perceptron (MLP, $d$-dimensional layers) with Gaussian Error Linear Unit (GELU) activation to further capture nonlinearities. This is connected to a residual layer to form the output $\mathbf{O}_{m}\in\mathbb{R}^{K\times L\times d}=\mathbf{O}_{r}+\texttt{MLP}(\mathbf{O}_{r})$. We then multiply $\mathbf{O}_{m}$ with a weight matrix $\mathbf{W}_{\text{out}}\in\mathbb{R}^{d\times d}$ to obtain the output of MVCA, $\mathbf{O}\in\mathbb{R}^{K\times L\times d}=\mathbf{O}_{m}\mathbf{W}_{\text{out}}$.
3.4 Koopman-guided spectrum evolvement
Motivated by Koopman operator theory [29, 40, 5], we introduce a layer to capture the temporal evolution of time-frequency patterns in wavelet space. We aim to learn a Koopman operator with $K$ dimensions in the transformed space. To obtain the operator $\mathbf{K}\in\mathbb{C}^{K\times K}$, we first initialise a learnable complex-valued matrix $\mathbf{S}\in\mathbb{C}^{K\times K}$. At each forward pass, we apply QR decomposition to $\mathbf{S}$, and retain only the resulting unitary matrix $\mathbf{U}\in\mathbb{C}^{K\times K}$, i.e. $\mathbf{U}\leftarrow\texttt{QR}\left(\mathbf{S}\right)$ s.t. $\mathbf{U}^{\dagger}\mathbf{U}=\mathbf{I}$, where $\mathbf{U}^{\dagger}$ is the conjugate transpose of $\mathbf{U}$. Multiplying with $\mathbf{U}$ therefore prevents data amplification or distortion.
We then initialise a learnable vector $\mathbf{p}\in\mathbb{R}^{K}$, where the $k$-th element, $p_{k}$, controls the temporal evolution (phase transformation angle) of the $k$-th transformation. All elements in the vector are then mapped into complex numbers, obtaining $\mathbf{v}\in\mathbb{C}^{K}$, where $v_{k}=e^{ip_{k}}$. We use $\mathbf{v}$ to form a diagonal matrix $\mathbf{D}\in\mathbb{C}^{K\times K}=\texttt{diag}(\mathbf{v})$. The Koopman operator, $\mathbf{K}$, is then given by $\mathbf{K}=\mathbf{U}\mathbf{D}\mathbf{U}^{\dagger}$. We use $\mathbf{K}$ to model the temporal evolution of the MVCA embedding, $\mathbf{O}$, after converting it to a complex form, $\mathbf{O}_{c}\in\mathbb{C}^{K\times L\times d}$ (imaginary part is set to $i0$):
$\mathbf{O}_{l}=\mathbf{K}\times\mathbf{O}_{c}\,\,,$ (4)
where $\mathbf{O}_{l}\in\mathbb{C}^{K\times L\times d}$ is the evolved embedding after multiplication with $\mathbf{K}$.
A conventional Koopman framework models temporal evolution recursively at each time step [28, 5]. We instead choose to apply Koopman transformation in the frequency domain with one forward pass, which is equivalent to learning a direct transformation from the input to the output. This global projection reduces the accumulation of sequential errors while maintaining robustness in training.
3.5 Sequence reconstruction from wavelets
The inverse transformation reconstructs the original sequence by aggregating weights from each wavelet atom. Given the evolved state $\mathbf{O}_{l}$, we first obtain its real part denoted as $\mathbf{O}_{r}\in\mathbb{R}^{K\times L\times d}$. For each wavelet atom indexed by $k$, let $\mathbf{O}_{k}\in\mathbb{R}^{L\times d}$ denote the corresponding slice of $\mathbf{O}_{r}$. We then multiply it with the wavelet atom $\mathbf{M}_{k}\in\mathbb{R}^{d\times L}$, i.e. $\mathbf{R}_{k}=\mathbf{O}_{k}\odot\mathbf{M}_{k}^{\top}$. $\mathbf{R}_{k}\in\mathbb{R}^{L\times d}$ denotes the series reconstruction from the $k$-th atom. This operation is conducted over all $K$ atoms. Joining all the reconstructed series forms a $K\times L\times d$ matrix. We sum over its first dimension to obtain the reconstructed embedding, denoted as $\mathbf{R}\in\mathbb{R}^{L\times d}$.
3.6 Convolutional decoder
Finally, a $3$-layer convolutional decoder transforms the learned representation. It comprises $3$ $1$-dimensional convolutional layers with GELU activations between every $2$ layers. Each layer uses kernel sizes $[5,3,3]$ and paddings $[2,1,1]$ respectively, followed by an adaptive average pooling layer at the end. The dimension of each layer is $[H\times 4,\,H\times 2,\,H]$, producing the final sequence representation $\mathbf{Z}_{\text{out}}\in\mathbb{R}^{H\times H}$, in accordance with the time steps of the target forecasting horizon. The result is then linearly projected using a weight vector $\mathbf{W}_{z}\in\mathbb{R}^{H}$ to generate the final output, as in $\hat{\mathbf{y}}\in\mathbb{R}^{H}=\mathbf{Z}_{\text{out}}\mathbf{W}_{z}$, in accordance with the dimensionality of the target variable.
4 Results
We assess forecasting accuracy using an expanded collection of data sets and tasks, to overcome potential biases present in the current literature. We first compare Sonnet against other competitive baseline models. We then investigate the role of attention mechanisms in time series forecasting models by removing or replacing naïve transformers with more advanced variants, including the proposed MVCA module. We also provide an ablation study of the key components of Sonnet.
4.1 Experiment settings
We conduct experiments over $12$ real-world data sets. This includes $2$ established benchmarks from prior papers [61, 58], specifically the ETTh1 and ETTh2 data sets, which contain hourly electricity transformer temperature forecasting. Oil temperature is our target variable, with the remaining indicators considered as exogenous variables. We also use $2$ data sets from the Darts library [16], predicting hourly energy prices (ENER) and electricity (low-voltage) consumption (ELEC). In addition, we form weather data sets extracted from the WeatherBench repository [38] for $5$ cities from diverse geographical locations—London (WEA-LD), New York (WEA-NY), Hong Kong (WEA-HK), Cape Town (WEA-CT), and Singapore (WEA-SG)—to support a more inclusive analysis. For each city, we sample data from its nearest grid point and obtain $5$ climate indicators. We provide spatial context by including data from its eight surrounding grid points ($3\times 3$ grid) as additional exogenous variables. Following prior work on global climate forecasting [49], we resample the data to a temporal resolution of $6$ hours. The forecasting target is the $850$ hPa (T850) temperature, a key indicator for climate modelling [41, 14]. Finally, we include influenza-like illness (ILI) rate forecasting tasks (as in [43]) in $3$ locations, England (ILI-ENG), and in U.S. Health & Human Services (HHS) Regions 2 (ILI-US2) and 9 (ILI-US9). In the ILI tasks, frequency time series of web searches are included as exogenous predictors. More information about the data sets is provided in Appendix A.
For the ETT tasks, we set the forecasting horizon ($H$) to $\{96,192,336,720\}$ time steps, and use a single test set of consecutive unseen instances adopting the evaluation settings in prior work [34, 25]. For ENER we set $H=\{24,48,72,168\}$ hours ahead and use $1$ year (2018) for testing. For ELEC, we use the last $2$ years (2020, ’21) as $2$ distinct test seasons, and set $H$ to $\{12,24,36\}$ hours. For the WEA tasks, we form $3$ test sets (years 2016, ’17, ’18), and set $H$ to $\{4,12,28,120\}$ time steps corresponding to $\{1,3,7,30\}$ days. Following the same setup as in [43], for the ILI forecasting task, we test models on $4$ consecutive influenza seasons (2015/16 to 2018/19), each time training a new model on data from previous seasons. We set $H=\{7,14,21,28\}$ days. Further details are provided in Appendix D.
We compare Sonnet to $8$ competitive forecasting models that, to the best of our knowledge, form the current SOTA methods: DLinear [58], Crossformer [60], iTransformer [25], PatchTST [34], TimeXer [52], Samformer [18], ModernTCN [27], and DeformTime [43]. We also include a naïve baseline which can either be a seasonal or a standard persistence model, depending on the presence of strong seasonality. For all tasks (except ETT), we conduct hyperparameter tuning for all models. For the ETT tasks, we adopt settings from the official repositories of methods (expect for Crossformer that did not provide this). Further details about the baseline models can be found in Appendix C.
Table 1: Forecasting accuracy results across all tasks, methods, and forecasting horizons ($H$). For the ELEC/ILI/WEA tasks, we report the average performance across $2$, $4$, and $3$ test sets, respectively (detailed breakdowns in Appendix E.1). $\epsilon\%$ denotes sMAPE. Best results are bolded and second best are underlined. Grey background denotes instances where a model does not outperform persistence.
Task Sonnet DeformTime ModernTCN Samformer TimeXer PatchTST iTransformer Crossformer DLinear
$H$ MAE $\epsilon\%$ MAE $\epsilon\%$ MAE $\epsilon\%$ MAE $\epsilon\%$ MAE $\epsilon\%$ MAE $\epsilon\%$ MAE $\epsilon\%$ MAE $\epsilon\%$ MAE $\epsilon\%$
ELEC 12 0.1040 24.95 0.1162 26.68 0.1596 36.81 0.2336 53.84 0.1287 28.45 0.1419 30.76 0.1468 31.91 0.1513 30.35 0.3307 71.51
24 0.1203 27.70 0.1359 30.11 0.1806 39.70 0.2520 56.39 0.1586 33.20 0.1545 34.04 0.1588 35.40 0.1873 35.60 0.3369 73.23
36 0.1389 30.21 0.1729 35.81 0.2065 43.79 0.2866 62.40 0.1785 35.33 0.1659 33.85 0.1791 38.18 0.2273 39.36 0.3908 82.38
ENER 24 0.3621 65.67 0.3717 66.92 0.3752 66.53 0.3703 66.01 0.3733 69.59 0.3717 65.89 0.3838 68.36 0.3716 67.44 0.4307 76.49
48 0.4120 71.07 0.4299 74.59 0.4154 71.74 0.4294 72.34 0.4404 77.61 0.4467 75.60 0.4386 75.63 0.4647 81.79 0.5138 89.79
72 0.4036 70.52 0.4217 74.50 0.4089 70.84 0.4353 73.40 0.4246 74.70 0.4303 73.90 0.4276 74.19 0.4473 78.62 0.5124 89.26
168 0.3980 68.63 0.4399 78.98 0.4497 74.86 0.4243 72.28 0.4493 76.76 0.4872 81.34 0.4352 74.11 0.4168 76.76 0.5070 87.87
ETTh1 96 0.2229 17.52 0.1941 14.96 0.2047 15.66 0.2047 15.73 0.2135 16.03 0.2017 15.41 0.2052 15.46 0.2126 16.52 0.2599 20.82
192 0.2414 18.62 0.2116 16.08 0.2417 18.32 0.2307 17.64 0.2322 16.96 0.2409 18.29 0.2429 18.13 0.2820 21.63 0.3798 31.78
336 0.2392 18.91 0.2158 16.27 0.2415 18.52 0.2523 18.94 0.2414 17.34 0.2559 19.29 0.2593 19.11 0.2947 22.65 0.6328 58.34
720 0.2989 24.28 0.2862 21.81 0.2785 20.44 0.3026 23.21 0.2617 18.58 0.3087 23.89 0.2886 22.05 0.3350 24.84 0.7563 69.52
ETTh2 96 0.3098 33.04 0.3121 40.07 0.3199 40.68 0.3312 40.16 0.3346 41.04 0.3145 39.25 0.3420 42.41 0.3486 40.71 0.3349 41.68
192 0.3742 39.58 0.3281 37.90 0.3887 47.08 0.3874 45.37 0.4154 47.07 0.3839 45.45 0.4233 47.44 0.4035 43.16 0.4084 50.67
336 0.3689 39.48 0.3450 37.00 0.3904 50.54 0.4083 46.41 0.4041 42.26 0.4018 46.77 0.4332 45.95 0.4487 49.44 0.4710 55.53
720 0.4335 51.62 0.3640 34.99 0.5728 63.04 0.5198 58.44 0.5135 56.17 0.4960 55.27 0.4565 45.40 0.5832 61.45 0.7981 94.67
ILI-ENG 7 1.4791 21.84 1.6417 28.60 1.9489 28.27 2.3475 28.31 2.8084 33.66 2.3115 27.61 2.3084 26.38 1.8698 25.70 2.8214 43.02
14 1.9225 25.77 2.2308 33.98 2.7050 36.01 3.0290 36.63 3.4937 41.88 3.2547 37.76 3.2301 36.67 2.6543 30.97 3.7922 55.29
21 2.5101 36.53 2.6500 32.70 3.0400 40.02 4.4980 54.41 4.3337 51.57 4.3192 51.11 4.2347 48.93 3.0014 40.57 4.4739 61.25
28 2.7481 36.95 2.7228 40.44 3.3611 47.87 5.1598 60.78 4.9013 61.60 4.9964 59.60 4.8125 55.35 3.1983 46.15 5.0347 67.75
ILI-US2 7 0.3806 14.89 0.4122 16.01 0.4398 16.55 0.6495 24.21 0.6083 23.38 0.7097 24.51 0.6507 23.24 0.4400 16.46 0.7355 27.94
14 0.4491 18.38 0.4752 17.73 0.5279 20.22 0.7696 30.16 0.7725 29.07 0.8635 30.11 0.7896 28.17 0.5852 20.98 0.8435 32.22
21 0.5326 20.64 0.5425 22.12 0.5781 23.85 0.8374 31.42 0.8243 31.46 1.0286 36.70 0.8042 30.03 0.6245 22.29 0.9124 34.93
28 0.5788 21.15 0.5538 22.25 0.5710 23.66 0.9389 36.80 0.9074 34.72 1.1525 42.61 0.9619 36.75 0.6512 23.47 0.9999 38.46
ILI-US9 7 0.2668 12.98 0.2622 12.26 0.2899 14.17 0.4025 19.39 0.3813 18.21 0.4116 19.34 0.4057 18.57 0.3149 14.44 0.4675 23.47
14 0.2806 13.10 0.3084 13.80 0.3417 15.29 0.5257 24.50 0.4665 22.14 0.5020 24.09 0.4702 22.44 0.3571 17.23 0.5467 27.35
21 0.3179 14.11 0.3179 14.23 0.3710 15.43 0.5415 24.22 0.5715 27.43 0.5935 29.40 0.5106 24.11 0.3418 15.90 0.6001 29.66
28 0.3675 16.53 0.3532 15.75 0.3940 17.19 0.6050 27.95 0.6555 31.32 0.6665 33.35 0.6498 31.05 0.3747 16.44 0.6564 32.16
WEA-CT 4 1.6240 9.63 1.7600 10.41 1.8752 11.05 2.2265 13.09 2.2376 13.14 3.5004 20.17 2.1906 12.83 1.6382 9.62 3.1483 18.22
12 3.5432 20.38 3.5681 20.46 3.7761 21.67 4.0444 23.14 3.7241 21.31 4.1910 23.99 3.9741 22.83 3.5932 20.62 4.0265 22.94
28 3.7277 21.32 3.7601 21.49 3.9399 22.36 3.9239 22.37 3.8325 21.87 3.9405 22.48 3.9584 22.60 3.8061 21.72 3.9254 22.37
120 3.7373 21.35 3.8040 21.67 4.0889 23.54 3.9018 22.30 3.8412 21.88 4.0547 23.02 3.9473 22.49 3.7659 21.49 4.0570 23.05
WEA-HK 4 0.6389 4.05 0.6804 4.32 0.7004 4.42 0.8097 5.10 0.8648 5.52 1.1488 7.15 0.8048 5.08 0.6488 4.11 0.9898 6.21
12 1.2355 7.74 1.2786 7.99 1.3555 8.43 1.4006 8.70 1.3027 8.13 1.5825 9.77 1.4225 8.86 1.2896 8.08 1.5464 9.62
28 1.4135 8.83 1.4746 9.16 1.5866 9.82 1.6226 10.05 1.5115 9.39 1.6441 10.16 1.6356 10.08 1.5282 9.45 1.6232 10.06
120 1.5469 9.58 1.5399 9.45 1.6326 10.22 1.8843 12.06 1.7092 10.55 2.0084 12.84 1.6745 10.45 1.5931 9.73 1.8329 11.49
WEA-LD 4 1.7231 15.16 1.8753 16.31 1.9456 16.88 2.1537 18.53 2.2628 19.25 2.7602 22.93 2.1509 18.55 1.7447 15.26 2.5065 20.94
12 2.9589 23.88 3.0214 24.15 3.2056 25.58 3.3070 26.61 3.1625 25.25 3.5406 28.33 3.3622 27.36 3.0492 24.35 3.3927 26.70
28 3.2161 25.49 3.2724 25.84 3.5067 27.48 3.5841 27.97 3.4672 27.23 3.7365 29.42 3.6884 29.14 3.3048 25.88 3.6073 28.16
120 3.2464 25.82 3.2973 26.17 3.8434 29.92 3.8420 30.32 3.6557 28.52 4.2344 32.40 3.8518 30.36 3.3935 26.75 3.9640 30.35
WEA-NY 4 1.2716 11.94 1.4028 13.03 1.4154 12.95 1.6003 14.44 1.7290 15.57 2.1644 19.24 1.6066 14.40 1.2935 12.24 1.9782 17.88
12 2.4476 21.23 2.4453 21.04 2.6221 22.93 2.7069 23.57 2.6537 22.85 2.8592 24.81 2.6609 23.09 2.4494 21.11 2.9507 25.29
28 2.6744 23.10 2.7450 23.20 2.9336 25.37 3.0347 26.18 2.8775 24.48 3.0956 27.01 3.0204 25.30 2.7830 23.66 3.2099 27.07
120 2.7135 23.15 2.8224 23.86 3.3000 28.05 3.5289 33.01 3.1501 26.55 3.4086 28.78 3.1029 27.74 2.9615 24.69 3.6129 29.75
WEA-SG 4 0.3444 1.25 0.3557 1.30 0.3624 1.32 0.3925 1.43 0.3801 1.38 0.4238 1.54 0.3868 1.41 0.3532 1.29 0.4048 1.47
12 0.4160 1.51 0.4256 1.55 0.4493 1.64 0.4784 1.74 0.4447 1.62 0.4992 1.82 0.4662 1.70 0.4359 1.59 0.4764 1.73
28 0.4653 1.69 0.4875 1.77 0.5196 1.89 0.5421 1.97 0.4974 1.81 0.5542 2.02 0.5307 1.93 0.5003 1.82 0.5284 1.92
120 0.4830 1.76 0.5050 1.83 0.5321 1.94 0.5221 1.90 0.5215 1.90 0.5357 1.95 0.5233 1.90 0.5212 1.89 0.5328 1.94
4.2 Forecasting accuracy of Sonnet
Prediction accuracy for all tasks is enumerated in Table 1, using Mean Absolute Error (MAE) and symmetric Mean Absolute Percentage Error (sMAPE or $\epsilon\%$) as the main evaluation metrics. For WEA tasks, sMAPE is modified to avoid small values due to Kelvin temperature units (see Appendix D.6). For the ETT and ENER tasks, results are based on a single forecasting season (i.e., one train / test split per forecasting horizon), while for other tasks we average over multiple seasons.
Sonnet exhibits the overall best performance across the explored MTS forecasting tasks. Comparing Sonnet’s performance to the best-performing baseline model for each task and forecasting horizon yields an MAE reduction of $1.1\%$ on average. Ranking-wise, Sonnet is the best-performing model on $34$, and the second-best on $9$ out of $47$ forecasting tasks. Sonnet shows relatively inferior performance on the ETT tasks. However, the ETT data sets have only $6$ exogenous variables and cover a limited training time span ($2$ years). Consequently, they offer a limited spectrum to gain from exploring variable dependencies, as well as insufficient historical context, both of which are elements that Sonnet leverages from. Aside from the ETT tasks, Sonnet reduces MAE by $3.3\%$ on average. Therefore, Sonnet is more effective with more covariates and longer time spans for training. Compared to a specific forecasting model, Sonnet consistently improves accuracy, reducing MAE from $2.2\%$ for the best-performing baseline (DeformTime) to $31.1\%$ for the worst-performing one (DLinear). This further highlights the model’s forecasting capacity amongst a wide range of applications.
Focusing on the more challenging forecasting tasks (ILI and WEA), Sonnet outperforms baselines $25$ or $26$ (out of $32$ tasks in total) based on sMAPE or MAE, respectively. The MAE reduction is $3.5\%$ on average for ILI tasks and $2\%$ for WEA tasks. Within these tasks, we also observe consistent comparative performance patterns in the baseline models. Models that do not capture inter-variable dependencies (DLinear and PatchTST) generally offer lower predictability (Sonnet reduces their MAE by $28.8\%$ and $29.8\%$), and in most ILI tasks, they cannot surpass the performance of a persistence model. In contrast, baseline models that capture inter-variable dependencies tend to perform better. Among these, those that embed the covariates along the temporal dimension—namely Samformer, TimeXer, and iTransformer—exhibit inferior performance (Sonnet reduces their MAE by $23.9\%$, $22.1\%$, and $22.9\%$), whereas models that preserve temporal order—namely DeformTime, ModernTCN, and Crossformer—achieve stronger results (Sonnet reduces their MAE by $3.4\%$, $11.1\%$, and $8\%$). Hence, in tasks with more informative covariates, preserving temporal structure while modelling inter-variable dependencies is a desirable property.
Table 2: Performance (average MAE across the $4$ test seasons) of iTransformer, Samformer, and PatchTST on the ILI forecasting tasks (ILI-ENG/US2/US9) with different modifications to the naïve attention mechanism. ‘$\neg$ Attn’ denotes the removal of the residual attention module, and FNet / FED / VDAB refer to using the attention modules proposed in FNet [43], FEDformer [62], and DeformTime [43], respectively. Best results are bolded and second best are underlined.
Attention $H=7$ days $H=14$ days $H=21$ days $H=28$ days
ENG US2 US9 ENG US2 US9 ENG US2 US9 ENG US2 US9
iTransformer 2.3084 0.6507 0.4057 3.2301 0.7896 0.4702 4.2347 0.8042 0.5106 4.8125 0.9619 0.6498
$\neg$ Attn 2.3011 0.7242 0.4475 3.2557 0.8696 0.5344 4.4729 1.0105 0.6207 5.1864 1.1667 0.7179
FNet 2.5534 0.7665 0.4741 3.5337 0.8746 0.5550 4.6970 1.0006 0.6066 5.3153 1.1118 0.7139
FED 3.6150 0.9286 0.5613 4.3098 1.0750 0.6527 5.8596 1.4140 0.8669 6.4196 1.5167 0.9397
VDAB 2.4287 0.6772 0.4126 3.3093 0.8274 0.4982 3.9406 0.8485 0.4992 4.4975 0.9068 0.5790
MVCA 2.2707 0.5483 0.3581 3.1233 0.7267 0.4454 4.1780 0.7578 0.4880 4.8705 0.8885 0.5409
SamFormer 2.3475 0.6495 0.4025 3.0290 0.7696 0.5257 4.4980 0.8374 0.5415 5.1598 0.9389 0.6050
$\neg$ Attn 2.5010 0.8245 0.5131 3.3015 0.9679 0.6047 4.1830 1.1396 0.7389 5.0136 1.2391 0.8347
FNet 2.4812 0.7765 0.4699 3.2054 0.9202 0.5865 4.3909 0.9266 0.6443 5.0592 1.0983 0.7149
FED 3.5217 0.9140 0.5599 4.2109 1.0519 0.6508 5.7800 1.3439 0.8519 5.9809 1.4348 0.9190
VDAB 2.0406 0.5755 0.3651 2.9525 0.7281 0.4505 4.1080 0.7883 0.4861 4.9015 0.8617 0.5605
MVCA 2.1365 0.5514 0.3609 3.2046 0.6935 0.4379 3.8033 0.7294 0.4975 4.7438 0.8159 0.5345
PatchTST 2.3115 0.7097 0.4116 3.2547 0.8635 0.5020 4.3192 1.0286 0.5935 4.9964 1.1525 0.6665
$\neg$ Attn 2.4723 0.7702 0.4585 3.6415 0.9017 0.5551 4.2998 1.0657 0.6472 4.8538 1.1704 0.7313
FNet 2.8097 0.8033 0.4591 4.0342 0.9322 0.5327 4.7975 1.0341 0.5430 4.9718 1.1087 0.6279
FED 3.6158 0.9292 0.5614 4.6572 1.0765 0.6542 5.8725 1.4043 0.8671 6.3728 1.4953 0.9306
VDAB 2.0799 0.5925 0.3820 3.0211 0.7722 0.4413 3.8164 0.8009 0.5159 4.6044 0.8518 0.5801
MVCA 2.0054 0.5824 0.3705 3.0411 0.7406 0.4291 3.8627 0.7871 0.4746 4.5712 0.8765 0.5491
4.3 Effectiveness of different attention modules in MTS forecasting
We evaluate the effectiveness of the proposed attention-driven module, MVCA, by integrating it into existing forecasting models (we refer to them as base models) that originally deploy naïve transformer attention. Experiments are conducted on the ILI tasks as these are the most representative of the MTS class we are exploring: they not only contain many exogenous predictors, but also frame hard practical epidemiological modelling problems when $H=21$ or $28$ days ahead. We use Samformer [18], iTransformer [25], and PatchTST [34] as base models. For each one of them, we evaluate $5$ attention configurations: removing the attention module altogether ($\neg$ Attn), attention with Fourier transformer proposed by FNet [21], the Frequency Enhanced Decomposed (FED) attention proposed in FEDformer [62], the variable deformable attention (VDAB) proposed by DeformTime [43], and ours (MVCA). Further details are provided in Appendices C.2 and D.8.
MAEs averaged across the $4$ test seasons for all forecasting horizons are enumerated in Table 2 (sMAPEs are reported in Appendix E.2, Table S9). Notably, for the base models, removing the attention module does not lead to significant performance degradation, with some cases even showing improved results, e.g. improved accuracy is reported for some horizons for the ILI-ENG task. This indicates that the attention modules in the base models do not always capture useful information.
By replacing naïve attention with the proposed MVCA module, we obtain significant performance gains across all ILI tasks, reducing MAE by $10.7\%$ on average across all base models. The average MAE reduction for PatchTST is greater ($15.1\%$) compared to the other base models ($10.4\%$ for Samformer and $6.7\%$ for iTransformer), indicating that the application of MVCA yields enhanced performance gains for the model that does not capture inter-variable dependencies. When compared to other modified attention modules, MVCA shows competitive performance, reducing MAE by $2.8\%$ on average. Compared to the best-performing baseline (VDAB from DeformTime), it reduces MAE and sMAPE by $3.5\%$ and $3\%$, respectively. While the gains in the ENG region are less pronounced, MVCA achieves consistently the best results across the US regions, and VDAB is the second best; both methods capture inter-variable dependencies while others do not. Comparably, replacing naïve attention with the Fourier attention modules from FEDformer and FNet results in worse performance.
4.4 Sonnet with different attention variants
To understand how different attention mechanisms affect the forecasting accuracy of Sonnet, we replace the MVCA module with the aforementioned modified attention mechanisms (see section 4.3). Experiments are conducted on both the ILI and WEA forecasting tasks across all locations, forecasting horizons, and test periods. Hyperparameters are re-tuned for each attention variant using the same validation procedure. Further experimental details are provided in Appendix C.2.
Table 3: Performance of Sonnet across the ILI and WEA tasks with different attention modules. Best results are bolded and second best are underlined.
Sonnet VDAB [43] FED [62] FNet [21] Naïve
$H$ MAE $\epsilon\%$ MAE $\epsilon\%$ MAE $\epsilon\%$ MAE $\epsilon\%$ MAE $\epsilon\%$
ILI-ENG 7 1.4791 21.84 1.4327 20.33 5.9913 77.67 1.5908 27.67 1.7170 25.34
14 1.9225 25.77 1.9729 29.39 6.0309 78.05 2.0841 32.81 2.0677 31.42
21 2.5101 36.53 2.7744 35.16 6.0447 78.24 2.7652 39.31 2.7812 37.17
28 2.7481 36.95 3.0602 44.31 6.0508 78.38 3.2403 44.51 3.0785 50.03
ILI-US2 7 0.3806 14.89 0.4294 16.57 1.1675 45.00 0.4146 18.50 0.4356 17.55
14 0.4491 18.38 0.4997 19.10 1.1792 45.53 0.5585 23.15 0.5316 20.43
21 0.5326 20.64 0.5748 20.49 1.1954 46.19 0.6593 25.31 0.6130 23.46
28 0.5788 21.15 0.6426 24.26 1.1934 46.13 0.7053 29.09 0.6514 26.12
ILI-US9 7 0.2668 12.98 0.2766 13.37 0.8558 40.97 0.2857 14.44 0.3006 13.85
14 0.2806 13.10 0.3239 15.81 0.8560 40.98 0.3510 16.13 0.3398 14.89
21 0.3179 14.11 0.3188 14.50 0.8567 41.01 0.3788 17.79 0.4038 19.23
28 0.3675 16.53 0.3748 16.81 0.8570 41.02 0.3791 18.20 0.4111 19.07
WEA-CT 4 1.6240 9.63 1.7135 10.11 4.3594 24.65 1.6287 9.63 1.7260 10.21
12 3.5432 20.38 3.5631 20.48 4.3595 24.65 3.5880 20.56 3.5581 20.47
28 3.7277 21.32 3.7602 21.50 4.3585 24.65 3.9155 22.32 3.7584 21.48
120 3.7373 21.35 3.7339 21.33 4.3564 24.64 3.8150 21.73 3.7632 21.50
Sonnet VDAB [43] FED [62] FNet [21] Naïve
$H$ MAE $\epsilon\%$ MAE $\epsilon\%$ MAE $\epsilon\%$ MAE $\epsilon\%$ MAE $\epsilon\%$
WEA-HK 4 0.6389 4.05 0.6451 4.11 2.8011 17.74 0.6389 4.08 0.6760 4.30
12 1.2355 7.74 1.2786 8.01 2.8008 17.74 1.3254 8.33 1.2810 8.06
28 1.4135 8.83 1.4633 9.06 2.8078 17.78 1.4941 9.30 1.4745 9.17
120 1.5469 9.58 1.5469 9.60 2.8100 17.80 1.6389 10.15 1.6211 10.04
WEA-LD 4 1.7231 15.16 1.8135 15.79 4.6682 34.93 1.7402 15.30 1.8107 15.99
12 2.9589 23.88 2.9863 24.06 4.6614 34.90 3.0078 24.11 2.9733 23.95
28 3.2161 25.49 3.2712 25.85 4.6593 34.87 3.3784 26.45 3.2682 25.84
120 3.2464 25.82 3.3193 26.24 4.6676 34.91 3.6666 28.48 3.4072 26.73
WEA-NY 4 1.2716 11.94 1.3307 12.36 5.3865 40.92 1.2684 11.77 1.3661 12.92
12 2.4476 21.23 2.4742 21.56 5.3836 40.91 2.5101 21.80 2.4380 21.45
28 2.6744 23.10 2.7081 23.43 5.3871 40.93 2.7972 23.81 2.7238 23.11
120 2.7135 23.15 2.7876 23.55 5.3892 40.93 3.0629 25.58 2.9106 24.55
WEA-SG 4 0.3444 1.25 0.3461 1.26 0.6999 2.54 0.3409 1.24 0.3455 1.26
12 0.4160 1.51 0.4207 1.53 0.6976 2.53 0.4238 1.54 0.4246 1.54
28 0.4653 1.69 0.4729 1.72 0.6979 2.53 0.4947 1.80 0.4712 1.71
120 0.4830 1.76 0.4957 1.80 0.6974 2.53 0.5119 1.86 0.4850 1.76
Results are enumerated in Table 3. Overall, using MVCA yields the best accuracy compared to other attention modules, outperforming them in $26$ out of the $32$ tasks (MAE). On average, it reduces the MAE and sMAPE scores by $2.9\%$ and $2\%$, respectively. The most competitive baseline module is VDAB (DeformTime). MVCA outperforms it by $3.6\%$ and $3\%$ on average w.r.t. MAE and sMAPE, respectively. The worst-performing attention module is FED (FEDformer). MVCA reduces its MAE by $51.5\%$. FED focuses on capturing seasonality information using frequency transformation without considering the dependencies between variables. This shows that modelling seasonality alone is insufficient for MTS forecasting, particularly when key predictive information comes from exogenous variables. The performance obtained by using either naïve attention or FNet is similar. MVCA reduces their MAE by $7\%$ and $6.6\%$ on average, respectively. This is an expected outcome as FNet does not introduce learnable parameters to the naïve attention.
Table 4: Ablation study results across the ILI (4 test seasons) and WEA tasks (3 test years). MAE is shown for each ablated variation ($\neg$ denotes removal of a component). $\Delta\%$ is the MAE percentage of difference compared to Sonnet’s MAE (which can be seen in Table 1).
$\neg$ Coherence $\neg$ MLP $\neg$ MVCA $\neg$ Embed $\neg$ Koopman
$H$ MAE $\Delta\%$ MAE $\Delta\%$ MAE $\Delta\%$ MAE $\Delta\%$ MAE $\Delta\%$
ILI-ENG 7 1.4224 –3.83 1.5312 +3.52 1.5906 +7.54 1.6333 +10.42 1.4855 +0.43
14 2.1099 +9.75 2.1710 +12.92 2.2346 +16.23 2.0658 +7.45 2.0793 +8.15
21 2.8797 +14.72 2.6141 +4.15 2.7939 +11.31 2.6467 +5.44 2.6299 +4.77
28 3.4200 +24.45 3.1145 +13.33 3.2745 +19.15 3.0910 +12.47 3.0622 +11.43
ILI-US2 7 0.4142 +8.82 0.4109 +7.96 0.4317 +13.43 0.4162 +9.37 0.4205 +10.48
14 0.5120 +14.00 0.5074 +12.96 0.5277 +17.50 0.5295 +17.90 0.4963 +10.51
21 0.6087 +14.29 0.5963 +11.96 0.6711 +26.02 0.5729 +7.57 0.5901 +10.81
28 0.6233 +7.67 0.6622 +14.39 0.6659 +15.05 0.6065 +4.78 0.6338 +9.50
ILI-US9 7 0.2720 +1.93 0.3017 +13.10 0.2775 +4.01 0.2784 +4.36 0.2796 +4.80
14 0.3183 +13.41 0.3341 +19.05 0.3307 +17.85 0.3146 +12.12 0.3151 +12.30
21 0.3575 +12.47 0.3264 +2.68 0.3929 +23.60 0.3113 –2.08 0.3192 +0.41
28 0.4277 +16.40 0.3446 –6.23 0.4013 +9.20 0.3587 –2.39 0.3476 –5.40
WEA-CT 4 1.6633 +2.42 1.6434 +1.20 1.6355 +0.71 1.6779 +3.32 1.6495 +1.57
12 3.5909 +1.35 3.5743 +0.88 3.5886 +1.28 3.5722 +0.82 3.5677 +0.69
28 3.7832 +1.49 3.7605 +0.88 3.8620 +3.60 3.7654 +1.01 3.7543 +0.72
120 3.7695 +0.86 3.7639 +0.71 3.7685 +0.83 3.7662 +0.77 3.7667 +0.79
$\neg$ Coherence $\neg$ MLP $\neg$ MVCA $\neg$ Embed $\neg$ Koopman
$H$ MAE $\Delta\%$ MAE $\Delta\%$ MAE $\Delta\%$ MAE $\Delta\%$ MAE $\Delta\%$
WEA-HK 4 0.6390 +0.00 0.6395 +0.08 0.6310 –1.24 0.6553 +2.55 0.6394 +0.07
12 1.2904 +4.45 1.2607 +2.04 1.3023 +5.41 1.2779 +3.43 1.2733 +3.07
28 1.4747 +4.33 1.4671 +3.80 1.4770 +4.50 1.4667 +3.76 1.4626 +3.48
120 1.6347 +5.68 1.6267 +5.16 1.6199 +4.72 1.5991 +3.37 1.6039 +3.68
WEA-LD 4 1.7538 +1.78 1.7675 +2.58 1.7420 +1.10 1.7868 +3.70 1.7555 +1.88
12 2.9906 +1.07 3.0070 +1.63 3.0009 +1.42 3.0187 +2.02 2.9926 +1.14
28 3.3126 +3.00 3.2812 +2.02 3.3278 +3.47 3.3261 +3.42 3.2681 +1.62
120 3.3854 +4.28 3.3468 +3.09 3.3749 +3.96 3.3591 +3.47 3.3454 +3.05
WEA-NY 4 1.2962 +1.94 1.2941 +1.77 1.2806 +0.71 1.3158 +3.48 1.2841 +0.99
12 2.4466 –0.04 2.4643 +0.68 2.4685 +0.85 2.4666 +0.78 2.4595 +0.49
28 2.7515 +2.88 2.7084 +1.27 2.7644 +3.36 2.7392 +2.42 2.7230 +1.82
120 2.8564 +5.27 2.7774 +2.36 2.9380 +8.27 2.8401 +4.67 2.8013 +3.24
WEA-SG 4 0.3417 –0.79 0.3446 +0.05 0.3426 –0.52 0.3481 +1.09 0.3431 –0.36
12 0.4226 +1.60 0.4194 +0.81 0.4260 +2.41 0.4228 +1.62 0.4220 +1.43
28 0.4894 +5.18 0.4755 +2.20 0.4742 +1.91 0.4754 +2.17 0.4759 +2.28
120 0.4999 +3.49 0.4942 +2.31 0.5028 +4.10 0.4965 +2.80 0.4950 +2.49
4.5 Ablation study on Sonnet
An ablation study is performed to better understand the contribution of Sonnet’s components to the overall forecasting accuracy. We ablate over the following modules: the coherence mechanism (within MVCA), the MLP head for inter-variable dependency modelling (within MVCA), the complete MVCA module (section 3.3), the joint embedding module (section 3.1), and the Koopman evolvement module (section 3.4). A more detailed description of each ablation variant is provided in Appendix D.9. We re-tune Sonnet’s hyperparameters for each ablation variant across all locations, test seasons, and forecasting horizons (for the ILI and WEA tasks).
Results are enumerated in Table 4. We report the average MAE across all test seasons and its percentage of difference compared to the Sonnet model (enumerated in Table 1). Among all the ablated components, the MVCA module has the greatest impact. Removing MVCA results in an increase of MAE by $6.3\%$ on average across the ILI and WEA tasks. Breaking down MVCA further, we observe that removing its two internal components—coherency modelling and the MLP residual connection—leads to notable drops in performance ($5.2\%$ and $4.1\%$ increase in MAE, respectively), but less pronounced than removing the entire MVCA block. This indicates that while capturing coherence over time and enhancing variable interactions with residual connected MLP are both important modules, they also work in tandem within the full MVCA design.
Other components within Sonnet also contribute to the model’s predictive ability, although to a lesser extent. Removing the joint embedding module leads to a $4\%$ increase in MAE, highlighting the importance of separately encoding the target and exogenous variables. Additionally, removing the Koopman operator module increases MAE by $3.2\%$, demonstrating the effectiveness of modelling the temporal evolution. Interestingly, as the forecasting horizon $H$ increases, the performance degradation becomes greater when removing coherence in MVCA, from an MAE increase of $2\%$ and $1.1\%$ (shortest $H$) for the ILI and WEA tasks, respectively, to $13.6\%$ and $3.7\%$ (longest $H$). This indicates that spectral coherence is more important for longer-term forecasting.
5 Conclusion
In this paper, we present Sonnet, a novel model for multivariable time series forecasting tasks. Sonnet operates in the spectral dimension using learnable wavelet transforms and captures variable dependencies with frequency-based coherence. It then predicts future temporal dynamics with the Koopman operator. Experiments are conducted on $12$ data sets from different application domains, with $9$ of them containing multiple test seasons for a more comprehensive analysis. Sonnet yields the best performance in $34$ out of $47$ tasks, reducing MAE by $2.2\%$ and $1.1\%$ on average compared to the most performant baseline overall (DeformTime) or per task (can vary), respectively. Additionally, our experiments highlight that replacing vanilla attention, present in various forecasting models, with the proposed MVCA module improves prediction accuracy by a large margin.
Acknowledgements
The authors would like to thank the RCGP for providing ILI rates for England. V. Lampos would like to acknowledge all levels of support from the EPSRC grant EP/X031276/1.
Appendix
Appendix A Data sets for multivariable time series forecasting
In this section, we provide a detailed description of all the data sets used in our MTS forecasting experiments. The weather data set that is formed by us is available for download at github.com/ClaudiaShu/Sonnet.
A.1 Hourly electricity transformer temperature prediction
The Hourly Electricity transformer temperature (ETTh1 and ETTh2) data sets are two widely used benchmarks in time series forecasting, obtained from $2$ different counties in China [61]. [^2] ETTh1 and ETTh2, github.com/zhouhaoyi/ETDataset The target variable of both data sets is the oil temperature. They contain $6$ exogenous variables capturing power load attributes. Both data sets cover the period from July 1, 2016, to June 26, 2018. We use only a subset of $14{,}400$ time steps starting from July 1, 2016, to match prior evaluation setups of the models that we are comparing to [34, 25, 43]. Within the subset, the first $8{,}640$ time steps are used for training, the next $2{,}880$ for validation, and the last $2{,}880$ for testing.
A.2 Electricity consumption prediction
The Electricity Consumption (ELEC) data set provides the electricity consumption records for Zurich (Switzerland), collected every $15$ minutes. [^3] Electricity consumption data set, github.com/unit8co/darts ELEC contains measurements for both residential households (recorded at low voltage) and commercial businesses (recorded at medium voltage). We choose the electricity consumption of households as our target variable for prediction. External predictors of the electricity consumption include $8$ weather-related covariates recorded hourly by a weather station in the city of Zurich.
Note that for some time intervals, certain predictor values may be missing. These are imputed using linear interpolation. We then resample rates to an hourly frequency following prior work in electricity consumption prediction [1, 57].
For the electricity consumption task, we use two test periods, years 2020 ($8{,}784$ time steps, including the leap day) and 2021 ($8{,}760$ time steps). The year preceding the test year is used as the validation ($8{,}760$ and $8{,}784$ time steps, respectively), while the $4$ years before validation form the training set ($35{,}064$ in both evaluations).
A.3 Energy price prediction
The energy price (ENER) data set contains $4$ years of electrical consumption, generation, and pricing data, with the main target of predicting the actual energy price in Spain. [^4] Energy price data set, github.com/unit8co/darts Note that the original data sets have $27$ exogenous variables in total. However, $8$ variables consist entirely of zeros or missing values, which do not provide useful information for forecasting purposes. We therefore exclude these variables from the analysis and use the remaining $19$. Data is sampled on an hourly basis. Data from December 31, 2014 to December 31, 2016 are used for training ($17{,}545$ samples). The subsequent year (January 1 to December 31, 2017, $8{,}760$ samples) is used for validation. For testing, we use data from January 1, 2018, to 22:00, December 31, 2018 ($8{,}759$ samples); the price for the last hour is not available.
A.4 Weather forecasting on selected locations
We form the weather forecasting (WEA) data set using climate information extracted from the WeatherBench repository [38], covering the period from 1979 to 2018.555WeatherBench repository, dataserv.ub.tum.de/index.php/s/m1524895 Data is sampled from a uniform spatial grid over the globe with a resolution of $5.625^{\circ}$ in both latitude and longitude. Each grid point represents a distinct geographical location. To capture different climatic conditions, we select five spatially diverse cities from the grid: London (UK), New York (US), Hong Kong (China), Cape Town (South Africa), and Singapore. Among these, London, New York, and Hong Kong are located in the northern hemisphere, with Hong Kong being in the subtropical zone; Singapore lies near the equator, and Cape Town is in the southern hemisphere. The two cities that are closer to the equator (Hong Kong and Singapore) do not have strong seasonal temperature fluctuations over the year. Consequently, we expect models to yield lower prediction errors at these two locations. For each city, climate data is sampled from the nearest grid point to its coordinates. [^6] We obtain the longitude and latitude of each city using OpenStreetMap (wiki.openstreetmap.org). Specifically, we obtain the absolute distance in both latitude and longitude between each city and all grid points, and select the grid point with the smallest distance along both axes. The forecasting target is the $850$ hPa temperature in Kelvin (T850) at the central grid point, [^7] Further information about T850 is provided at charts.ecmwf.int/products/medium-z500-t850. which is an important climate indicator that has been widely used in climate research [41, 14]. To form exogenous inputs, we include four other climate variables at each selected grid point following prior work [33, 49]. These are the $500$ hPa geopotential (Z500), $2$-meter air temperature (t2m), $10$-meter zonal wind (u10), and $10$-meter meridional wind (v10). We also include these $4$ variables and T850 from the surrounding eight grid points, forming a $3\times 3$ spatial window centred on each selected grid, resulting in $44$ exogenous variables in total. The data is resampled to maintain a temporal resolution of $6$ hours similarly to Verma et al. [49].
We consider $3$ test seasons: years 2016, 2017, and 2018 (Jan. 1 to Dec. 31). Each test year contains $1{,}464$, $1{,}460$, and $1{,}460$ time steps, respectively. The year immediately preceding a test year is used for validation (2015, 2016, and 2017). Validation year 2015 contains $1{,}460$ time steps. The training sets include data from January 1, 1980, up to the date preceding the corresponding validation year, resulting in $51{,}136$, $52{,}596$, and $54{,}060$ time steps for the three test seasons.
A.5 Influenza-like-illness rate prediction
We construct an ILI forecasting data set for three regions: England (UK), and US HHS Regions 2 and 9, following the setup in prior work [43]. Ground truth ILI rates are obtained from the Royal College of General Practitioners (RCGP) in England and the Centers for Disease Control and Prevention (CDC) for the US regions. Weekly ILI rates are linearly interpolated to a daily resolution. Specifically, we assume the ILI rates are reported corresponding to the middle day of each week, which is Wednesday for the US regions and Thursday for England [43]. We also incorporate the ILI rate reporting delay of syndromic surveillance systems; in England there is a lag of $\delta=7$ days and $\delta=14$ days for the US regions [50, 39].
We use web search activity data (daily frequency time series of different search terms), obtained from the Google Health Trends API, as the exogenous indicators. We follow the setup of DeformTime [43] to form the set of search queries for our experiments. [^8] Search queries can be obtained from github.com/ClaudiaShu/DeformTime. Note that for the US regions, search frequencies are collected at the state level, and the query data for each region is obtained using weighted averages based on the population of the states inside each region (New Jersey/New York is $0.32/0.68$ for Region 2 and Arizona/California is $0.16/0.84$ for Region 9).
We evaluate forecasting accuracy using $4$ test seasons (2015/16, 2016/17, 2017/18, and 2018/19) in England and the US regions, each corresponding to a 12-month period that encompasses at least one seasonal flu outbreak. For England, each season begins on September $1$ and ends on August $31$ of the next year, while for the US regions, it is from August $1$ to July $31$ of the following year. For each test season, models are trained on the preceding $9$ seasons. Validation sets are extracted from the last three seasons in the training data. Each validation set covers $180$ days, containing three $60$-day segments that capture the onset (from the last season), peak (from the penultimate season), and outset (the $3$rd to last season) phases of an influenza season. The onset is defined as the first point followed by two weeks above a predefined ILI threshold; the peak is the point of the highest ILI rate; and the outset is the last point followed by two weeks above the threshold. These selected points are then considered as the $30$th day in a $60$-day validation period.
Appendix B Further information about MVCA
This section supplements section 3.3 of the main paper. Here, we further clarify complex number operations.
In Equation 2, the cross-spectral and power-spectral density matrices are obtained using the frequency domain embeddings $\mathbf{Q}_{f}$, $\mathbf{K}_{f}\in\mathbb{C}^{L\times\ell}$. The obtained matrices $\mathbf{P}_{qq}=\mathbf{Q}_{f}\odot\mathbf{Q}_{f}^{\ast}$ and $\mathbf{P}_{kk}=\mathbf{K}_{f}\odot\mathbf{K}_{f}^{\ast}$ are both containing real numbers because multiplying a complex number (e.g. $z=a+ib$) by its conjugate ($a-ib$) results in its squared magnitude ($a^{2}+b^{2}$), which is a real number.
In Equation 3, we obtain the coherence vector $\mathbf{C}_{qk}$ using the averaged cross-spectral density matrix $\mathbf{P}_{qk}\in\mathbf{C}^{L\times\ell}$. The average of this imaginary matrix $\mathbf{P}_{qk}$ is obtained by averaging the real and imaginary components separately. Taking the magnitude $|\overline{\mathbf{P}}_{qk}|$ then results in a real-valued vector, as magnitude is essentially the Euclidean norm (i.e., the square root of the sum of the squares of the real and imaginary parts) of each complex entry. As such, we obtain a real-valued coherence vector $\mathbf{C}_{qk}$.
Appendix C Baseline forecasting models and attention mechanisms
In this section, we first provide a brief description of the baseline models that we use in our experiments (section C.1). To the best of our knowledge, they form the current state-of-the-art (SOTA) in MTS forecasting. These also include the base models that use naïve attention mechanisms, which we then replace with MVCA (section 4.3). We also describe the attention modules that have been used in related literature and which we compare to MVCA (section C.2).
C.1 Baseline MTS forecasting models
DeformTime [43] is specifically designed for multivariable forecasting. It uses deformable attention to capture the dependencies across time and covariates, reaching competitive performance on MTS forecasting tasks. Based on our evaluation, DeformTime is the best-performing baseline.
ModernTCN [27] applies convolution along the temporal and variable axes. By leveraging large receptive fields, it effectively captures both intra-variable and inter-variable dependencies.
TimeXer [52] is a transformer-based model that focuses on multivariable forecasting. It first embeds the exogenous variables along the sequence dimension, and then captures the exogenous information through a cross-attention mechanism.
Samformer [18] first embeds the input along the sequence dimension and then applies naïve attention along the feature dimension. It conducts forecasts using a linear layer. We also include Samformer as one of the base models that deploy naïve attention.
iTransformer [25] embeds the input along the sequence dimension and deploys multiple attention layers along the variable dimension. We also include iTransformer as one of the base models.
PatchTST [34] segments time series data into patches of fixed size, treating them as input tokens to a transformer. It processes each variable independently, enabling efficient modelling of long-term dependencies. PatchTST is also used as one of the base models.
Crossformer [60] captures both inter-variable and intra-variable dependencies using a two-stage attention mechanism, with each stage dedicated to one type of dependency. The information is captured using time series patches.
DLinear [58] uses a simpler neural network architecture (feedforward layers). It decomposes each input series into trend and seasonal components, which are then forecasted independently using linear layers. It models each variable separately, using its own historical values for prediction. Based on our evaluation, DLinear is the worst-performing model in MTS forecasting. It does not outperform persistence in the majority of our experiments (see Table 1). This corroborates findings in related work [43].
Naïve forecasting models. We include simple baseline models as reference points for performance comparison. We use either seasonal persistence or standard persistence, depending on whether a target variable (and task) is characterised by strong seasonality or not. A standard persistence model is a naïve baseline that uses the last observed value of the target variable as the prediction (forecasting outcome). The seasonal persistence model uses the value from the same point in the previous seasonal cycle. In our case, the electricity consumption data (ELEC) has a strong seasonal trend, with seasonal persistence outperforming the standard one by a large margin (results not shown). Therefore, we use the seasonal persistence model for the ELEC task, while for the other tasks, we retain standard persistence.
Table S1: ILI rate forecasting results in England (ILI-ENG) with web search frequency time series as exogenous variables. We assess performance on $4$ test seasons (2015/16 to 2018/19) using linear correlation ($r$), MAE, and sMAPE ($\epsilon\%$). Best results are bolded and second best are underlined.
$H$ Model 2016 2017 2018 2019 Average
$r$ MAE $\epsilon\%$ $r$ MAE $\epsilon\%$ $r$ MAE $\epsilon\%$ $r$ MAE $\epsilon\%$ $r$ MAE $\epsilon\%$
7 DLinear 0.8710 2.6985 38.86 0.8769 1.6879 33.77 0.8231 4.2734 50.14 0.8906 2.6258 49.30 0.8654 2.8214 43.02
iTransformer 0.8773 2.0645 27.26 0.8911 1.3734 24.06 0.8180 4.1077 31.26 0.8911 1.6881 22.94 0.8694 2.3084 26.38
PatchTST 0.8832 1.9823 26.13 0.8744 1.4307 25.39 0.8475 3.7925 31.23 0.8526 2.0403 27.70 0.8644 2.3115 27.61
TimeXer 0.8927 1.9807 26.29 0.8773 1.5327 29.94 0.6974 5.1695 43.18 0.7564 2.5507 35.22 0.8060 2.8084 33.66
Samformer 0.8780 2.0244 27.60 0.8922 1.4025 25.10 0.7718 4.1902 35.01 0.8857 1.7729 25.55 0.8569 2.3475 28.31
ModernTCN 0.9490 1.6043 17.82 0.9606 1.1988 22.75 0.9679 3.2340 40.55 0.9316 1.7584 31.98 0.9523 1.9489 28.27
Crossformer 0.8879 2.3314 36.08 0.9591 0.8606 17.82 0.9234 2.9133 27.04 0.9436 1.3737 21.89 0.9285 1.8698 25.70
DeformTime 0.9728 1.3786 19.57 0.9665 1.3083 29.49 0.9585 2.6313 41.40 0.9436 1.2485 23.96 0.9603 1.6417 28.60
Sonnet 0.9621 1.0294 13.44 0.9772 1.2003 20.79 0.9590 2.0369 25.10 0.9590 1.6498 28.02 0.9643 1.4791 21.84
14 DLinear 0.7810 3.6354 52.29 0.7857 2.1872 42.47 0.6834 5.5368 60.91 0.8351 3.8092 65.47 0.7713 3.7922 55.29
iTransformer 0.7731 2.7620 36.46 0.7776 1.9929 34.17 0.6576 5.7796 44.34 0.7850 2.3862 31.71 0.7483 3.2301 36.67
PatchTST 0.8078 2.6267 35.59 0.7730 1.9562 33.55 0.6722 5.6075 43.64 0.7218 2.8284 38.25 0.7437 3.2547 37.76
TimeXer 0.7891 2.6580 35.27 0.7910 1.9864 36.99 0.5503 6.2678 53.64 0.6646 3.0626 41.61 0.6988 3.4937 41.88
Samformer 0.7205 3.0697 41.49 0.8113 1.7285 28.83 0.6888 4.9426 39.40 0.8246 2.3754 36.78 0.7613 3.0290 36.63
ModernTCN 0.8814 2.1727 31.93 0.8864 1.4953 28.61 0.8905 4.6759 48.74 0.8441 2.4760 34.76 0.8756 2.7050 36.01
Crossformer 0.7940 2.9124 41.75 0.8843 1.4711 23.39 0.8470 3.8295 30.88 0.9007 2.4044 27.86 0.8565 2.6543 30.97
DeformTime 0.9259 2.0556 25.36 0.9400 1.3180 33.86 0.8964 3.7631 47.33 0.9154 1.7863 29.36 0.9194 2.2308 33.98
Sonnet 0.9025 1.5327 22.20 0.9512 1.2453 21.28 0.9233 2.8313 28.62 0.9307 2.0808 30.99 0.9269 1.9225 25.77
21 DLinear 0.4466 4.1958 56.02 0.6441 2.8294 52.56 0.4920 6.9514 71.66 0.6608 3.9192 64.75 0.5609 4.4739 61.25
iTransformer 0.5310 4.1390 55.62 0.5987 2.6718 43.65 0.4403 7.0816 55.17 0.6556 3.0464 41.29 0.5564 4.2347 48.93
PatchTST 0.6126 3.6607 48.57 0.6214 2.6377 44.77 0.4427 7.3828 60.54 0.5861 3.5956 50.55 0.5657 4.3192 51.11
TimeXer 0.5958 3.7654 49.46 0.6981 2.4045 41.17 0.4377 7.3686 62.71 0.5301 3.7962 52.91 0.5654 4.3337 51.57
Samformer 0.4919 4.2610 57.57 0.5352 2.8996 43.99 0.5263 6.8716 60.01 0.4793 3.9598 56.07 0.5082 4.4980 54.41
ModernTCN 0.7869 2.3283 27.42 0.8590 1.6790 30.86 0.9085 5.7489 61.11 0.8357 2.4040 40.67 0.8475 3.0400 40.02
Crossformer 0.8113 2.6710 36.42 0.8014 1.8035 31.29 0.8504 4.5615 49.08 0.7331 2.9697 45.48 0.7991 3.0014 40.57
DeformTime 0.8819 2.1700 27.01 0.8859 1.8980 32.65 0.9110 4.8984 44.78 0.9092 1.6335 26.36 0.8970 2.6500 32.70
Sonnet 0.8537 1.9394 25.77 0.8720 1.5679 29.12 0.8260 3.9711 43.60 0.9042 2.5620 47.62 0.8640 2.5101 36.53
28 DLinear 0.3917 4.5039 60.30 0.5078 3.3928 60.94 0.3911 7.7738 77.58 0.5774 4.4683 72.19 0.4670 5.0347 67.75
iTransformer 0.3492 4.9542 65.45 0.6761 2.2745 38.38 0.2947 8.3463 69.39 0.5286 3.6750 48.18 0.4622 4.8125 55.35
PatchTST 0.4271 4.5240 60.22 0.4477 3.2658 53.01 0.3203 8.0693 65.91 0.4486 4.1265 59.27 0.4109 4.9964 59.60
TimeXer 0.4978 4.2217 56.56 0.5523 2.8584 48.91 0.3413 8.1493 71.44 0.3337 4.3758 69.50 0.4313 4.9013 61.60
Samformer 0.3995 4.7015 62.97 0.3194 3.5264 55.69 0.3127 8.1404 66.57 0.3636 4.2709 57.89 0.3488 5.1598 60.78
ModernTCN 0.8017 2.6486 36.17 0.7628 2.6508 51.35 0.7544 4.8228 49.19 0.8152 3.3222 54.78 0.7835 3.3611 47.87
Crossformer 0.7299 3.1253 41.56 0.8082 1.8197 33.57 0.7841 5.4882 55.97 0.9349 2.3599 53.48 0.8143 3.1983 46.15
DeformTime 0.7919 2.6540 49.29 0.9229 1.8100 41.87 0.8493 4.8324 37.53 0.9499 1.5947 33.06 0.8785 2.7228 40.44
Sonnet 0.8975 2.0133 31.59 0.8841 1.7490 33.24 0.7745 4.4605 41.57 0.7935 2.7698 41.41 0.8374 2.7481 36.95
C.2 Attention mechanisms used for time series forecasting
Variable Deformable Attention Block (VDAB). Introduced in DeformTime [43], VDAB adaptively attends to relevant information across time and variables by applying cross attention over the original and deformed input. This enables to capture complex inter-variable dependencies in MTS forecasting.
Frequency Enhanced Decomposition (FED). Employed in FEDformer [62], FED decomposes time series into trend and seasonal components, applying attention mechanisms to each, thereby capturing both periodic patterns and time-delay dependencies effectively. We note that we do not compare to the complete forecasting model of FEDformer as it has been outperformed by most of our selected baseline models. Crossformer, iTransformer, PatchTST, ModernTCN, DLinear, and Samformer have all provided direct comparison with FEDformer in their original papers. Meanwhile, the ones that do not provide a direct comparison with FEDformer, namely TimeXer and DeformTime, have both outperformed PatchTST and iTransformer on MTS forecasting tasks by a large margin. Given Sonnet’s competitive performance over all the aforementioned (baseline) models, we expect it to yield superior performance compared to FEDformer.
FNet. The attention proposed in FNet [21] replaces the commonly used self-attention mechanism in a transformer with a parameter-free Fourier transform. Specifically, it maps the input sequence to the frequency domain and retains only the real part of the result. This approach enables effective mixing between input tokens along the time dimension without adding learnable parameters. FNet is not included as a baseline model in the main results as it is not a model designed for forecasting purposes. Additionally, we show in sections 4.3 and 4.4 that it gives a similar performance to the vanilla transformer, which has been outperformed.
Naïve Attention. The commonly used attention module was first proposed by Vaswani et al. [48] as a core component of the transformer model. The attention weights are computed using scaled dot product operations between query and key embeddings, facilitating the model’s ability to focus on relevant parts of the input sequence.
Table S2: ILI rate forecasting results in US Region 2 (ILI-US2) with web search frequency time series as exogenous variables. We assess performance on $4$ test seasons (2015/16 to 2018/19) using linear correlation ($r$), MAE, and sMAPE ($\epsilon\%$). Best results are bolded and second best are underlined.
$H$ Model 2016 2017 2018 2019 Average
$r$ MAE $\epsilon\%$ $r$ MAE $\epsilon\%$ $r$ MAE $\epsilon\%$ $r$ MAE $\epsilon\%$ $r$ MAE $\epsilon\%$
7 DLinear 0.7456 0.4982 29.23 0.7980 0.8210 31.85 0.7133 1.0910 30.82 0.8955 0.5318 19.84 0.7881 0.7355 27.94
iTransformer 0.7986 0.3865 21.30 0.8171 0.7367 27.07 0.7784 0.9080 23.22 0.8545 0.5717 21.38 0.8122 0.6507 23.24
PatchTST 0.7666 0.4290 23.78 0.7973 0.7546 26.97 0.6937 1.1018 26.52 0.8624 0.5533 20.79 0.7800 0.7097 24.51
TimeXer 0.8024 0.4214 24.76 0.8480 0.6882 27.16 0.8494 0.8233 22.90 0.8859 0.5001 18.70 0.8464 0.6083 23.38
Samformer 0.7549 0.4346 23.83 0.7935 0.7784 28.57 0.8563 0.8725 25.30 0.8834 0.5126 19.16 0.8220 0.6495 24.21
ModernTCN 0.8745 0.3874 21.99 0.9359 0.5216 17.45 0.9569 0.4782 14.12 0.9535 0.3722 12.64 0.9302 0.4398 16.55
Crossformer 0.8796 0.3382 21.39 0.9107 0.4975 18.32 0.9009 0.6309 15.61 0.9523 0.2936 10.52 0.9109 0.4400 16.46
DeformTime 0.8887 0.3428 21.86 0.9463 0.4796 18.66 0.9008 0.5369 12.88 0.9622 0.2894 10.64 0.9245 0.4122 16.01
Sonnet 0.9027 0.2854 16.40 0.9336 0.5471 18.68 0.9683 0.3708 11.52 0.9537 0.3192 12.97 0.9396 0.3806 14.89
14 DLinear 0.6546 0.5761 33.62 0.7335 0.9406 36.90 0.6007 1.2213 34.62 0.8261 0.6360 23.74 0.7037 0.8435 32.22
iTransformer 0.6864 0.4894 26.35 0.7548 0.8934 33.26 0.6025 1.1639 30.28 0.8294 0.6115 22.78 0.7183 0.7896 28.17
PatchTST 0.6628 0.5103 27.27 0.7350 0.9267 34.44 0.5993 1.3085 31.82 0.7922 0.7085 26.90 0.6973 0.8635 30.11
TimeXer 0.7178 0.4823 27.24 0.7446 0.9211 36.47 0.7118 1.0505 28.09 0.8164 0.6362 24.47 0.7476 0.7725 29.07
Samformer 0.7204 0.4929 28.49 0.7359 0.9580 37.34 0.8198 0.9533 29.98 0.7851 0.6741 24.81 0.7653 0.7696 30.16
ModernTCN 0.8093 0.4647 25.78 0.9000 0.6482 23.18 0.9360 0.6020 17.05 0.9572 0.3968 14.87 0.9006 0.5279 20.22
Crossformer 0.7872 0.4357 24.61 0.8640 0.6502 22.95 0.8057 0.8159 20.83 0.9241 0.4389 15.54 0.8453 0.5852 20.98
DeformTime 0.8050 0.4404 23.82 0.9271 0.5029 18.53 0.9126 0.6351 16.34 0.9652 0.3226 12.21 0.9025 0.4752 17.73
Sonnet 0.8654 0.3477 20.08 0.8832 0.6254 21.81 0.9731 0.4380 15.30 0.9499 0.3854 16.34 0.9179 0.4491 18.38
21 DLinear 0.6016 0.6023 35.11 0.6897 0.9671 38.06 0.4966 1.3272 38.76 0.7936 0.7530 27.80 0.6454 0.9124 34.93
iTransformer 0.6418 0.4790 24.54 0.7515 0.8784 34.22 0.8230 1.0845 31.85 0.7321 0.7748 29.52 0.7371 0.8042 30.03
PatchTST 0.5775 0.5592 29.41 0.6286 1.1661 43.41 0.5189 1.4395 38.13 0.6478 0.9495 35.86 0.5932 1.0286 36.70
TimeXer 0.4856 0.5815 32.17 0.7739 0.8696 33.97 0.6529 1.1238 32.12 0.7482 0.7225 27.56 0.6651 0.8243 31.46
Samformer 0.6042 0.5472 28.91 0.7109 0.9739 35.70 0.7263 1.1426 35.51 0.7586 0.6860 25.58 0.7000 0.8374 31.42
ModernTCN 0.8104 0.4656 26.17 0.8746 0.6790 30.30 0.9105 0.7258 20.75 0.9487 0.4421 18.19 0.8860 0.5781 23.85
Crossformer 0.7729 0.4634 26.54 0.8996 0.6688 25.62 0.7621 0.8882 23.17 0.9354 0.4777 13.82 0.8425 0.6245 22.29
DeformTime 0.7414 0.4568 25.94 0.9028 0.6189 26.44 0.9313 0.6981 21.04 0.9408 0.3963 15.09 0.8791 0.5425 22.12
Sonnet 0.7670 0.4571 25.22 0.9164 0.6012 23.16 0.8722 0.6961 18.91 0.9536 0.3759 15.28 0.8773 0.5326 20.64
28 DLinear 0.4836 0.6512 37.96 0.6112 1.0328 41.19 0.4125 1.4045 40.27 0.6840 0.8337 31.08 0.5479 0.9805 37.63
iTransformer 0.5258 0.5949 32.54 0.6567 1.0632 41.38 0.5726 1.3053 38.82 0.6738 0.8842 34.26 0.6072 0.9619 36.75
PatchTST 0.4013 0.6678 36.55 0.5274 1.2615 48.85 0.3903 1.6234 44.67 0.5579 1.0576 40.39 0.4692 1.1525 42.61
TimeXer 0.3979 0.5856 32.82 0.7017 0.9748 39.33 0.6137 1.2462 35.29 0.7037 0.8229 31.45 0.6042 0.9074 34.72
Samformer 0.4934 0.5858 31.55 0.6947 1.0083 39.01 0.6215 1.4670 50.33 0.7698 0.6945 26.31 0.6448 0.9389 36.80
ModernTCN 0.7635 0.4702 27.09 0.8667 0.6742 27.31 0.9223 0.6618 19.83 0.9604 0.4776 20.41 0.8782 0.5710 23.66
Crossformer 0.7102 0.4697 27.38 0.9064 0.7624 28.93 0.8432 0.8253 20.69 0.8782 0.5474 18.64 0.8345 0.6512 23.91
DeformTime 0.8369 0.4716 27.21 0.9562 0.6017 25.66 0.9253 0.5954 17.62 0.9294 0.5463 18.52 0.9119 0.5538 22.25
Sonnet 0.7578 0.4614 25.04 0.9140 0.6458 22.49 0.8747 0.7363 21.05 0.9609 0.4717 16.01 0.8768 0.5788 21.15
Appendix D Supplementary experiment settings
This section supplements section 4.1 in the main paper. All experiments were conducted using a Linux server (Ubuntu 24.04.2) with 3 NVIDIA L40S GPUs, 2 AMD EPYC 9354 CPUs, and 768GB of DDR5 RAM, except for the ablation experiments which were conducted using a Linux server (Ubuntu 22.04.5) with 2 NVIDIA A40 GPUs, 2 AMD EPYC 7443 CPUs, and 512GB of DDR4 RAM.
D.1 Hyperparameter settings
The look-back window $L$ is set differently depending on the forecasting task. For the ETTh1, ETTh2 data sets, $L$ is set to $336$ time steps for all $4$ different forecasting horizons $H=\{96,192,336,720\}$, based on prior work [34]. In ELEC experiments, we set $L=168$ hours (a week, given the strong weekly periodicity in these time series) for all $H=\{12,24,36\}$. For ENER, $L$ is set equal to the forecasting horizon, i.e. $L=H=\{24,48,72,168\}$. For the WEA tasks, we use a look-back window $L=\{28,28,56,240\}$ time steps for $H=\{4,12,28,120\}$, correspondingly. This means that we use the past $7$ days to conduct short-term forecasts ($1$ and $3$ days ahead), and the past $2H$ days for longer-term forecasts. For the ILI tasks, we set $L=\{28,28,56,56\}$ days respectively for forecasting horizons $H=\{7,14,21,28\}$ days ahead following prior work [43].
The batch size $d$ is set to $64$ for all tasks except for the ETT ones, where it is a learnable hyperparameter (see section D.4). We train using a maximum of $100$ epochs (see our early stopping settings in  D.4). Neural networks are optimised with Adam using mean squared error (MSE) loss on all time steps. Note that the output can be a time series (sequence) of one variable (the target) or more variables (target and exogenous), depending on the forecasting method.
Table S3: ILI rate forecasting results in US Region 9 (ILI-US9) with web search frequency time series as exogenous variables. We assess performance on $4$ test seasons (2015/16 to 2018/19) using linear correlation ($r$), MAE, and sMAPE ($\epsilon\%$). Best results are bolded and second best are underlined.
$H$ Model 2016 2017 2018 2019 Average
$r$ MAE $\epsilon\%$ $r$ MAE $\epsilon\%$ $r$ MAE $\epsilon\%$ $r$ MAE $\epsilon\%$ $r$ MAE $\epsilon\%$
7 DLinear 0.8071 0.4769 25.07 0.8092 0.3727 22.02 0.6759 0.5704 22.96 0.8925 0.4499 23.82 0.7962 0.4675 23.47
iTransformer 0.8380 0.3736 17.55 0.8446 0.2721 16.22 0.6664 0.6856 26.00 0.9230 0.2915 14.52 0.8180 0.4057 18.57
PatchTST 0.8248 0.3993 19.30 0.8318 0.2802 16.91 0.7068 0.6078 23.25 0.8796 0.3592 17.91 0.8107 0.4116 19.34
TimeXer 0.8237 0.4071 20.04 0.8471 0.2669 15.50 0.7638 0.5235 20.92 0.9021 0.3277 16.36 0.8342 0.3813 18.21
Samformer 0.8086 0.4116 19.72 0.8045 0.3049 18.05 0.7880 0.5042 19.90 0.8663 0.3893 19.91 0.8169 0.4025 19.39
ModernTCN 0.9263 0.2615 13.08 0.8886 0.2482 14.86 0.8706 0.4745 20.37 0.9770 0.1754 8.37 0.9156 0.2899 14.17
Crossformer 0.9305 0.2556 12.21 0.9524 0.2874 14.89 0.8213 0.4524 17.08 0.9435 0.2640 13.58 0.9119 0.3149 14.44
DeformTime 0.9161 0.2437 12.46 0.9356 0.2364 11.69 0.8744 0.3664 13.61 0.9675 0.2023 11.29 0.9234 0.2622 12.26
Sonnet 0.9301 0.2668 12.94 0.9293 0.2692 16.42 0.8481 0.3376 11.89 0.9693 0.1935 10.67 0.9192 0.2668 12.98
14 DLinear 0.7323 0.5767 29.99 0.7371 0.4384 25.44 0.6105 0.6635 27.69 0.8216 0.5085 26.27 0.7254 0.5467 27.35
iTransformer 0.7837 0.4310 20.63 0.7629 0.3448 20.49 0.6138 0.6853 27.21 0.8495 0.4196 21.41 0.7525 0.4702 22.44
PatchTST 0.7626 0.4810 23.33 0.7411 0.3524 21.48 0.6246 0.7164 28.11 0.8102 0.4583 23.44 0.7346 0.5020 24.09
TimeXer 0.7423 0.4795 23.26 0.7872 0.3078 18.07 0.6443 0.6440 25.49 0.8261 0.4345 21.76 0.7500 0.4665 22.14
Samformer 0.7297 0.4972 23.71 0.7555 0.3416 20.14 0.6032 0.7957 30.46 0.8117 0.4683 23.68 0.7250 0.5257 24.50
ModernTCN 0.9250 0.3582 15.44 0.8543 0.3175 17.00 0.8696 0.4244 15.32 0.9320 0.2666 13.40 0.8952 0.3417 15.29
Crossformer 0.8787 0.3333 16.05 0.8965 0.2980 16.74 0.8339 0.4325 16.91 0.9169 0.3647 19.21 0.8815 0.3571 17.23
DeformTime 0.9033 0.2962 14.39 0.9374 0.2893 13.15 0.8490 0.3897 14.62 0.9378 0.2584 13.05 0.9069 0.3084 13.80
Sonnet 0.9244 0.2807 12.91 0.9049 0.2832 14.71 0.8771 0.3490 13.24 0.9615 0.2096 11.51 0.9170 0.2806 13.10
21 DLinear 0.6687 0.6032 30.66 0.6920 0.4960 28.49 0.5830 0.7252 30.15 0.7778 0.5760 29.33 0.6804 0.6001 29.66
iTransformer 0.8659 0.3645 17.75 0.7078 0.3702 21.73 0.4561 0.8642 33.94 0.8298 0.4433 23.01 0.7149 0.5106 24.11
PatchTST 0.6899 0.5705 28.41 0.6399 0.4255 26.14 0.5214 0.7963 32.80 0.7137 0.5815 30.27 0.6412 0.5935 29.40
TimeXer 0.6203 0.5899 29.37 0.6373 0.4132 24.19 0.5158 0.7825 31.10 0.7460 0.5006 25.04 0.6298 0.5715 27.43
Samformer 0.7079 0.5086 24.07 0.8118 0.3403 18.93 0.5685 0.7794 27.50 0.7169 0.5378 26.37 0.7013 0.5415 24.22
ModernTCN 0.8695 0.3118 12.39 0.8775 0.3695 16.88 0.7572 0.5057 18.36 0.8997 0.2968 14.08 0.8510 0.3710 15.43
Crossformer 0.9565 0.3267 15.49 0.8669 0.3168 15.95 0.8777 0.3964 15.11 0.9135 0.3273 17.07 0.9036 0.3418 15.90
DeformTime 0.8331 0.3543 15.62 0.8966 0.3039 14.99 0.9141 0.3536 13.06 0.9249 0.2598 13.27 0.8922 0.3179 14.23
Sonnet 0.8573 0.2888 13.12 0.9619 0.3270 14.18 0.8359 0.4013 14.37 0.9382 0.2545 14.75 0.8983 0.3179 14.11
28 DLinear 0.6083 0.6496 32.81 0.6016 0.5329 30.18 0.5506 0.7893 32.75 0.7438 0.6539 32.89 0.6261 0.6564 32.16
iTransformer 0.6984 0.5531 27.80 0.5779 0.4444 25.70 0.3211 0.9942 40.08 0.6797 0.6077 30.60 0.5693 0.6498 31.05
PatchTST 0.6067 0.6457 32.75 0.5192 0.4903 30.18 0.4424 0.8809 36.70 0.6555 0.6492 33.78 0.5560 0.6665 33.35
TimeXer 0.5345 0.6372 32.39 0.6320 0.4302 23.80 0.3998 0.9296 37.46 0.6127 0.6248 31.64 0.5448 0.6555 31.32
Samformer 0.6543 0.5652 27.78 0.6646 0.3909 22.79 0.4984 0.8367 30.81 0.5980 0.6273 30.40 0.6038 0.6050 27.95
ModernTCN 0.8552 0.3611 17.38 0.9076 0.3748 17.59 0.8068 0.5156 18.99 0.8645 0.3245 14.79 0.8585 0.3940 17.19
Crossformer 0.8969 0.3398 16.47 0.8346 0.3647 16.99 0.8514 0.4365 15.42 0.8923 0.3578 16.87 0.8688 0.3747 16.44
DeformTime 0.8888 0.3718 15.55 0.9046 0.3153 14.57 0.9037 0.4185 16.18 0.9260 0.3069 16.68 0.9058 0.3532 15.75
Sonnet 0.8947 0.3072 15.45 0.9288 0.3723 16.09 0.8850 0.4497 16.70 0.8904 0.3406 17.90 0.8997 0.3675 16.53
D.2 Hyperparameter settings specific to Sonnet
[…]
D.3 Hyperparameter settings specific to the ILI rate forecasting tasks
[…]
D.4 Hyperparameter validation
[…]
Table S4: WEA forecasting results (temperature at $850$ hPa) in London (WEA-LD). We assess performance on $3$ test periods (years 2016, ’17, ’18) using linear correlation ($r$), MAE, and sMAPE ($\epsilon\%$). Best results are bolded and second best are underlined.
$H$ Model 2016 2017 2018 Average
$r$ MAE $\epsilon\%$ $r$ MAE $\epsilon\%$ $r$ MAE $\epsilon\%$ $r$ MAE $\epsilon\%$
4 DLinear 0.8341 2.3772 15.26 0.7730 2.5122 16.47 0.8422 2.6301 31.10 0.8164 2.5065 20.94
iTransformer 0.8746 2.0626 13.39 0.8325 2.1816 14.51 0.8803 2.2087 27.76 0.8625 2.1509 18.55
PatchTST 0.7975 2.6606 17.02 0.7233 2.7843 18.32 0.8077 2.8359 33.44 0.7762 2.7602 22.93
TimeXer 0.8707 2.1636 14.07 0.8256 2.2316 14.69 0.8681 2.3932 29.00 0.8548 2.2628 19.25
Samformer 0.8705 2.0649 13.40 0.8393 2.1271 14.15 0.8752 2.2692 28.04 0.8617 2.1537 18.53
ModernTCN 0.8972 1.8388 11.88 0.8560 2.0076 13.33 0.9063 1.9904 25.41 0.8865 1.9456 16.88
Crossformer 0.9186 1.6316 10.67 0.8788 1.8383 12.14 0.9225 1.7641 22.97 0.9066 1.7447 15.26
DeformTime 0.9060 1.7647 11.64 0.8620 1.9632 12.78 0.9096 1.8982 24.50 0.8926 1.8753 16.31
Sonnet 0.9206 1.6571 10.95 0.8887 1.7880 11.85 0.9269 1.7242 22.67 0.9121 1.7231 15.16
12 DLinear 0.6875 3.2620 20.14 0.5777 3.2740 20.77 0.6782 3.6422 39.20 0.6478 3.3927 26.70
iTransformer 0.7073 3.1987 20.42 0.6167 3.2760 21.05 0.6916 3.6120 40.61 0.6719 3.3622 27.36
PatchTST 0.6801 3.4350 21.80 0.5729 3.4499 22.13 0.6639 3.7370 41.06 0.6390 3.5406 28.33
TimeXer 0.7286 3.0578 19.05 0.6602 3.0519 19.57 0.7360 3.3778 37.11 0.7082 3.1625 25.25
Samformer 0.7117 3.1711 20.14 0.6352 3.1860 20.37 0.6969 3.5638 39.31 0.6812 3.3070 26.61
ModernTCN 0.7266 3.0716 19.39 0.6476 3.1372 19.95 0.7140 3.4080 37.39 0.6961 3.2056 25.58
Crossformer 0.7530 2.9360 18.52 0.6326 3.1054 19.77 0.7626 3.1061 34.75 0.7161 3.0492 24.35
DeformTime 0.7403 2.9613 18.51 0.6810 2.9529 18.83 0.7600 3.1499 35.10 0.7271 3.0214 24.15
Sonnet 0.7648 2.8561 17.98 0.6937 2.9154 18.69 0.7652 3.1051 34.98 0.7413 2.9589 23.88
28 DLinear 0.6197 3.5890 21.92 0.5479 3.3473 21.14 0.6560 3.8855 41.41 0.6078 3.6073 28.16
iTransformer 0.6016 3.7784 23.47 0.5710 3.4687 22.17 0.6506 3.8182 41.77 0.6077 3.6884 29.14
PatchTST 0.6130 3.7670 23.35 0.5358 3.6196 23.20 0.6436 3.8229 41.71 0.5974 3.7365 29.42
TimeXer 0.6561 3.5155 21.56 0.6287 3.2048 20.43 0.6842 3.6813 39.69 0.6563 3.4672 27.23
Samformer 0.6239 3.6455 22.48 0.5648 3.4061 21.46 0.6594 3.7008 39.99 0.6160 3.5841 27.97
ModernTCN 0.6509 3.4012 20.79 0.5680 3.4846 22.16 0.6841 3.6345 39.48 0.6343 3.5067 27.48
Crossformer 0.6791 3.2933 20.29 0.6278 3.1441 19.89 0.7001 3.4770 37.46 0.6690 3.3048 25.88
DeformTime 0.7076 3.2300 19.96 0.6311 3.1675 20.08 0.7298 3.4197 37.48 0.6895 3.2724 25.84
Sonnet 0.7171 3.0995 19.21 0.5982 3.2051 20.27 0.7307 3.3438 37.00 0.6820 3.2161 25.49
120 DLinear 0.5178 4.0071 24.30 0.3307 3.7275 23.27 0.6050 4.1572 43.48 0.4845 3.9640 30.35
iTransformer 0.5382 3.9632 23.80 0.5504 3.4798 22.72 0.5855 4.1123 44.57 0.5581 3.8518 30.36
PatchTST 0.3849 4.5359 27.91 0.3942 3.9432 25.12 0.5567 4.2240 44.16 0.4453 4.2344 32.40
TimeXer 0.6451 3.4743 21.42 0.4839 3.5218 22.15 0.6744 3.9711 42.00 0.6011 3.6557 28.52
Samformer 0.5761 3.8839 24.10 0.4873 3.6875 23.55 0.6481 3.9545 43.31 0.5705 3.8420 30.32
ModernTCN 0.5646 3.7511 22.77 0.5422 3.5645 23.40 0.5976 4.2146 43.60 0.5682 3.8434 29.92
Crossformer 0.6996 3.3054 20.39 0.6396 3.0820 19.72 0.6691 3.7933 40.13 0.6694 3.3935 26.75
DeformTime 0.7277 3.1270 19.63 0.6141 3.1808 20.31 0.7058 3.5842 38.58 0.6826 3.2973 26.17
Sonnet 0.7281 3.1081 19.38 0.6365 3.1281 20.03 0.6899 3.5029 38.04 0.6849 3.2464 25.82
D.5 Data normalisation
We apply z-score normalisation to the data of all forecasting tasks. This standardises endogenous and exogenous variables in a training set to have a zero mean and unit standard deviation. Normalisation is transferred from the training data to the test data. Data is reverted back to the original scale when computing prediction errors. Additionally, for the ETT, ENER, and ELEC data sets, we adopt variable normalisation [19] over the target variable within the input’s look-back window following prior work [34, 43].
Table S5: WEA forecasting results (temperature at $850$ hPa) in New York (WEA-NY). We assess performance on $3$ test periods (years 2016, ’17, ’18) using linear correlation ($r$), MAE, and sMAPE ($\epsilon\%$). Best results are bolded and second best are underlined.
$H$ Model 2016 2017 2018 Average
$r$ MAE $\epsilon\%$ $r$ MAE $\epsilon\%$ $r$ MAE $\epsilon\%$ $r$ MAE $\epsilon\%$
4 DLinear 0.9285 1.9481 16.49 0.9182 2.1349 21.52 0.9320 1.8516 15.65 0.9262 1.9782 17.88
iTransformer 0.9476 1.5907 12.82 0.9412 1.7162 17.87 0.9488 1.5129 12.53 0.9459 1.6066 14.40
PatchTST 0.9038 2.1592 17.66 0.8927 2.3817 23.79 0.9148 1.9524 16.26 0.9038 2.1644 19.24
TimeXer 0.9402 1.7540 14.49 0.9382 1.8287 18.59 0.9460 1.6045 13.63 0.9415 1.7290 15.57
Samformer 0.9478 1.5838 13.06 0.9415 1.7441 17.92 0.9520 1.4729 12.35 0.9471 1.6003 14.44
ModernTCN 0.9607 1.3723 11.37 0.9539 1.5332 15.92 0.9614 1.3408 11.57 0.9587 1.4154 12.95
Crossformer 0.9676 1.2417 10.31 0.9631 1.3846 15.29 0.9687 1.2541 11.14 0.9665 1.2935 12.24
DeformTime 0.9580 1.4031 11.30 0.9592 1.4651 15.92 0.9638 1.3403 11.85 0.9603 1.4028 13.03
Sonnet 0.9677 1.2506 10.43 0.9645 1.3370 14.70 0.9691 1.2271 10.69 0.9671 1.2716 11.94
12 DLinear 0.8464 2.8374 22.86 0.8060 3.1841 29.99 0.8487 2.8307 23.01 0.8337 2.9507 25.29
iTransformer 0.8626 2.5160 20.10 0.8275 2.9724 29.06 0.8638 2.4942 20.12 0.8513 2.6609 23.09
PatchTST 0.8405 2.7056 21.91 0.7976 3.2096 30.96 0.8471 2.6623 21.56 0.8284 2.8592 24.81
TimeXer 0.8686 2.4893 19.91 0.8404 2.9183 27.85 0.8744 2.5534 20.79 0.8611 2.6537 22.85
Samformer 0.8646 2.5356 20.66 0.8293 2.9936 29.05 0.8594 2.5914 20.98 0.8511 2.7069 23.57
ModernTCN 0.8677 2.4940 20.69 0.8332 2.9126 28.23 0.8689 2.4598 19.87 0.8566 2.6221 22.93
Crossformer 0.8893 2.3005 18.45 0.8570 2.7562 26.36 0.8916 2.2916 18.52 0.8793 2.4494 21.11
DeformTime 0.8852 2.3050 18.55 0.8519 2.7148 26.08 0.8864 2.3162 18.48 0.8745 2.4453 21.04
Sonnet 0.8858 2.3262 18.74 0.8499 2.7186 26.25 0.8902 2.2980 18.71 0.8753 2.4476 21.23
28 DLinear 0.8127 3.1132 24.71 0.7533 3.4967 32.30 0.8217 3.0198 24.19 0.7959 3.2099 27.07
iTransformer 0.8227 2.8172 21.76 0.7710 3.5240 32.26 0.8387 2.7201 21.89 0.8108 3.0204 25.30
PatchTST 0.8167 2.9402 23.82 0.7579 3.4486 33.28 0.8256 2.8982 23.92 0.8000 3.0956 27.01
TimeXer 0.8333 2.8010 22.43 0.8052 3.2027 30.03 0.8541 2.6288 20.97 0.8309 2.8775 24.48
Samformer 0.8221 2.8725 23.07 0.7663 3.3700 32.19 0.8285 2.8616 23.29 0.8056 3.0347 26.18
ModernTCN 0.8365 2.7179 21.53 0.7673 3.3990 32.94 0.8474 2.6840 21.66 0.8170 2.9336 25.37
Crossformer 0.8602 2.5428 20.23 0.8107 3.1558 29.80 0.8563 2.6502 20.96 0.8424 2.7830 23.66
DeformTime 0.8636 2.5246 19.87 0.8166 3.0717 28.77 0.8548 2.6388 20.96 0.8450 2.7450 23.20
Sonnet 0.8545 2.5533 20.23 0.8342 2.9812 29.08 0.8719 2.4887 19.99 0.8535 2.6744 23.10
120 DLinear 0.7883 3.4169 26.88 0.6870 3.9611 34.99 0.7596 3.4606 27.39 0.7450 3.6129 29.75
iTransformer 0.8418 2.7846 22.62 0.7691 3.5547 37.41 0.8123 2.9695 23.20 0.8078 3.1029 27.74
PatchTST 0.7964 3.1199 24.02 0.7451 3.6650 34.57 0.7857 3.4409 27.76 0.7757 3.4086 28.78
TimeXer 0.8198 2.9056 23.44 0.8003 3.2102 30.24 0.7830 3.3346 25.97 0.8010 3.1501 26.55
Samformer 0.7795 3.3766 28.57 0.7268 3.9559 42.16 0.7942 3.2543 28.30 0.7668 3.5289 33.01
ModernTCN 0.7883 3.1432 24.11 0.7353 3.5156 34.01 0.7781 3.2413 26.02 0.7672 3.3000 28.05
Crossformer 0.8382 2.7770 21.74 0.8143 3.1789 29.76 0.8129 2.9286 22.59 0.8218 2.9615 24.69
DeformTime 0.8642 2.5570 20.65 0.8013 3.2431 30.09 0.8529 2.6672 20.85 0.8395 2.8224 23.86
Sonnet 0.8614 2.5606 20.76 0.8377 2.9614 28.01 0.8597 2.6185 20.68 0.8529 2.7135 23.15
D.6 sMAPE definition for the weather (temperature) prediction tasks
We note that for the WEA tasks, the unit of the temperature time series sequence is originally captured in Kelvin.99900 degrees in Celsius is equal to $273.15$ Kelvin. Directly computing the sMAPE using Kelvin units will lead to a relatively small error (MAEs are not affected, but sMAPEs will be significantly lower and harder to display in an accessible way). To address this, we subtract the minimum value of the target variable across the test set, and then add a constant number $a=30$ to both the ground truth and the predicted value to avoid division by 00. This alters the formula of sMAPE as follows:
$\text{sMAPE}(\mathbf{y},\hat{\mathbf{y}})=\frac{100}{L}\sum_{t=1}^{L}\frac{|y_{t}-\hat{y}_{t}|}{\left(|y_{t}-\min(\mathbf{y})+a|+|\hat{y}_{t}-\min(\mathbf{y})+a|\right)/2}\times 100\%\,\,.$ (S1)
D.7 Random seed initialisation
We use a fixed seed equal to ‘42’ during training for all forecasting tasks, except for the ETTh1 and ETTh2 data sets. For these data sets, we follow prior work [34, 58] and use ‘2021’, with the exception of iTransformer, which uses ‘2023’ in line with its official configuration.
Table S6: WEA forecasting results (temperature at $850$ hPa) in Hong Kong (WEA-HK). We assess performance on $3$ test periods (years 2016, ’17, ’18) using linear correlation ($r$), MAE, and sMAPE ($\epsilon\%$). Best results are bolded and second best are underlined.
$H$ Model 2016 2017 2018 Average
$r$ MAE $\epsilon\%$ $r$ MAE $\epsilon\%$ $r$ MAE $\epsilon\%$ $r$ MAE $\epsilon\%$
4 DLinear 0.8955 1.0544 7.13 0.9249 0.9475 5.29 0.9091 0.9675 6.21 0.9098 0.9898 6.21
iTransformer 0.9343 0.8432 5.80 0.9462 0.7692 4.25 0.9404 0.8020 5.20 0.9403 0.8048 5.08
PatchTST 0.8489 1.2284 8.28 0.8955 1.0634 5.87 0.8612 1.1546 7.31 0.8686 1.1488 7.15
TimeXer 0.9350 0.9353 6.53 0.9452 0.8258 4.63 0.9515 0.8333 5.41 0.9439 0.8648 5.52
Samformer 0.9340 0.8454 5.82 0.9455 0.7852 4.32 0.9414 0.7984 5.15 0.9403 0.8097 5.10
ModernTCN 0.9549 0.7132 4.93 0.9589 0.6858 3.77 0.9551 0.7023 4.57 0.9563 0.7004 4.42
Crossformer 0.9573 0.6755 4.67 0.9654 0.6245 3.46 0.9628 0.6463 4.19 0.9618 0.6488 4.11
DeformTime 0.9554 0.7276 5.03 0.9626 0.6519 3.61 0.9620 0.6617 4.31 0.9600 0.6804 4.32
Sonnet 0.9617 0.6666 4.63 0.9677 0.6253 3.45 0.9673 0.6249 4.08 0.9656 0.6389 4.05
12 DLinear 0.7259 1.6528 11.11 0.8036 1.4281 7.90 0.7237 1.5581 9.85 0.7511 1.5464 9.62
iTransformer 0.7632 1.5487 10.49 0.8319 1.3230 7.29 0.7864 1.3958 8.80 0.7938 1.4225 8.86
PatchTST 0.7110 1.6836 11.32 0.7943 1.4387 7.87 0.6995 1.6251 10.11 0.7349 1.5825 9.77
TimeXer 0.7916 1.4408 9.70 0.8623 1.2177 6.75 0.8406 1.2495 7.94 0.8315 1.3027 8.13
Samformer 0.7804 1.5051 10.18 0.8340 1.3189 7.28 0.7891 1.3778 8.65 0.8012 1.4006 8.70
ModernTCN 0.7819 1.4543 9.81 0.8482 1.2875 7.12 0.8026 1.3248 8.35 0.8109 1.3555 8.43
Crossformer 0.7613 1.4767 9.95 0.8716 1.1834 6.57 0.8463 1.2087 7.73 0.8264 1.2896 8.08
DeformTime 0.8084 1.4218 9.50 0.8780 1.1443 6.38 0.8351 1.2698 8.09 0.8405 1.2786 7.99
Sonnet 0.8114 1.3921 9.35 0.8820 1.1220 6.25 0.8548 1.1923 7.62 0.8494 1.2355 7.74
28 DLinear 0.6844 1.6762 11.32 0.7547 1.5545 8.56 0.6757 1.6389 10.30 0.7049 1.6232 10.06
iTransformer 0.6986 1.6872 11.39 0.7499 1.6095 8.79 0.6914 1.6101 10.07 0.7133 1.6356 10.08
PatchTST 0.6935 1.6950 11.50 0.7515 1.5721 8.64 0.6758 1.6653 10.34 0.7069 1.6441 10.16
TimeXer 0.7459 1.5904 10.71 0.8013 1.4330 7.92 0.7581 1.5111 9.55 0.7684 1.5115 9.39
Samformer 0.6904 1.6976 11.49 0.7735 1.5271 8.43 0.6872 1.6431 10.25 0.7170 1.6226 10.05
ModernTCN 0.6928 1.7091 11.51 0.7954 1.4720 8.06 0.7131 1.5787 9.89 0.7338 1.5866 9.82
Crossformer 0.6899 1.7047 11.32 0.8231 1.4059 7.75 0.7642 1.4739 9.29 0.7590 1.5282 9.45
DeformTime 0.7721 1.5560 10.36 0.8288 1.3821 7.67 0.7724 1.4858 9.44 0.7911 1.4746 9.16
Sonnet 0.7714 1.5099 10.11 0.8518 1.2950 7.22 0.7718 1.4355 9.16 0.7983 1.4135 8.83
120 DLinear 0.6432 1.9457 13.25 0.6875 1.8548 10.28 0.7085 1.6983 10.94 0.6797 1.8329 11.49
iTransformer 0.7564 1.6063 11.01 0.7150 1.6955 9.34 0.6740 1.7218 10.98 0.7152 1.6745 10.45
PatchTST 0.6846 2.1676 15.78 0.6216 2.0684 11.43 0.7092 1.7893 11.30 0.6718 2.0084 12.84
TimeXer 0.7208 1.7303 11.64 0.6869 1.7013 9.33 0.7211 1.6961 10.69 0.7096 1.7092 10.55
Samformer 0.6898 1.9304 13.80 0.6868 1.8838 10.46 0.7217 1.8388 11.93 0.6995 1.8843 12.06
ModernTCN 0.7470 1.6455 11.27 0.7985 1.5017 8.31 0.6635 1.7506 11.09 0.7363 1.6326 10.22
Crossformer 0.6789 1.7296 11.34 0.7945 1.4771 8.09 0.7373 1.5724 9.77 0.7369 1.5931 9.73
DeformTime 0.7630 1.6101 10.65 0.7953 1.5073 8.23 0.7459 1.5023 9.46 0.7681 1.5399 9.45
Sonnet 0.7563 1.7449 11.65 0.8106 1.4349 7.92 0.7739 1.4610 9.17 0.7803 1.5469 9.58
D.8 Replacing attention modules with MVCA
To use MVCA in place of vanilla attention in the $3$ base models for experiments in section 4.3, some modifications are needed as MVCA initially expects a $3$-dimensional input in $\mathbb{R}^{K\times L\times d}$ with $K$ as the number of wavelet transformations (which we use as the number of heads in multi-head MVCA). In contrast, the attention module in the three base models takes input embeddings with only $2$ dimensions $\in\mathbb{R}^{u\times v}$, which represent time and variables. To address this discrepancy, we split the embedding into multiple attention heads, as MVCA is essentially a variant similar to multihead attention. Specifically, given an embedding in shape $\mathbb{R}^{u\times v}$, we first split it along the feature dimension into $K$ heads, resulting in an input dimension of $v/K$ for each head, where the head embeddings are in shape $\mathbb{R}^{K\times u\times(v/K)}$. The remaining operations then follow the standard MVCA process, with $u$ and $v/K$ corresponding to $L$ and $d$ respectively. The output of different heads is also joined along the feature dimension to form the final output in shape $\mathbb{R}^{u\times v}$. In our experiment, the number of heads is fixed to $8$ for all base models when replacing the attention with MVCA.
Table S7: WEA forecasting results (temperature at $850$ hPa) in Singapore (WEA-SG). We assess performance on $3$ test periods (years 2016, ’17, ’18) using linear correlation ($r$), MAE, and sMAPE ($\epsilon\%$). Best results are bolded and second best are underlined.
$H$ Model 2016 2017 2018 Average
$r$ MAE $\epsilon\%$ $r$ MAE $\epsilon\%$ $r$ MAE $\epsilon\%$ $r$ MAE $\epsilon\%$
4 DLinear 0.7760 0.4068 1.47 0.7495 0.4174 1.52 0.7885 0.3903 1.43 0.7713 0.4048 1.47
iTransformer 0.8033 0.3835 1.38 0.7777 0.3985 1.45 0.8089 0.3783 1.39 0.7967 0.3868 1.41
PatchTST 0.7572 0.4194 1.51 0.7266 0.4364 1.59 0.7578 0.4156 1.53 0.7472 0.4238 1.54
TimeXer 0.7933 0.3910 1.41 0.8010 0.3780 1.38 0.8147 0.3712 1.36 0.8030 0.3801 1.38
Samformer 0.7917 0.3940 1.42 0.7690 0.4064 1.48 0.8109 0.3770 1.38 0.7905 0.3925 1.43
ModernTCN 0.8194 0.3664 1.32 0.8135 0.3662 1.34 0.8286 0.3546 1.30 0.8205 0.3624 1.32
Crossformer 0.8198 0.3695 1.33 0.8403 0.3441 1.26 0.8407 0.3462 1.27 0.8336 0.3532 1.29
DeformTime 0.8166 0.3737 1.35 0.8356 0.3452 1.26 0.8349 0.3481 1.28 0.8290 0.3557 1.30
Sonnet 0.8313 0.3564 1.29 0.8363 0.3431 1.25 0.8469 0.3336 1.22 0.8382 0.3444 1.25
12 DLinear 0.6873 0.4789 1.72 0.6565 0.4844 1.77 0.6639 0.4658 1.71 0.6693 0.4764 1.73
iTransformer 0.7134 0.4571 1.65 0.6811 0.4714 1.72 0.6759 0.4702 1.72 0.6901 0.4662 1.70
PatchTST 0.6632 0.4880 1.76 0.6201 0.5099 1.86 0.6236 0.4995 1.83 0.6356 0.4992 1.82
TimeXer 0.7178 0.4516 1.63 0.7203 0.4404 1.61 0.7084 0.4421 1.62 0.7155 0.4447 1.62
Samformer 0.6976 0.4696 1.69 0.6613 0.4875 1.78 0.6680 0.4779 1.75 0.6756 0.4784 1.74
ModernTCN 0.7328 0.4458 1.61 0.7179 0.4445 1.62 0.7022 0.4576 1.68 0.7177 0.4493 1.64
Crossformer 0.7326 0.4548 1.64 0.7615 0.4124 1.50 0.7283 0.4406 1.61 0.7408 0.4359 1.59
DeformTime 0.7336 0.4319 1.56 0.7551 0.4203 1.53 0.7325 0.4246 1.56 0.7404 0.4256 1.55
Sonnet 0.7456 0.4302 1.55 0.7752 0.4009 1.46 0.7437 0.4169 1.53 0.7548 0.4160 1.51
28 DLinear 0.6046 0.5229 1.88 0.5531 0.5403 1.97 0.5513 0.5220 1.92 0.5697 0.5284 1.92
iTransformer 0.5986 0.5261 1.90 0.5899 0.5339 1.95 0.5652 0.5322 1.95 0.5846 0.5307 1.93
PatchTST 0.5614 0.5468 1.97 0.5414 0.5593 2.04 0.5210 0.5566 2.04 0.5413 0.5542 2.02
TimeXer 0.6816 0.4966 1.79 0.6231 0.5027 1.83 0.6152 0.4930 1.81 0.6399 0.4974 1.81
Samformer 0.5995 0.5262 1.90 0.5566 0.5498 2.00 0.5384 0.5503 2.02 0.5648 0.5421 1.97
ModernTCN 0.6151 0.5183 1.87 0.6013 0.5186 1.89 0.5871 0.5220 1.92 0.6012 0.5196 1.89
Crossformer 0.6660 0.5212 1.88 0.6690 0.5088 1.85 0.6692 0.4711 1.73 0.6681 0.5003 1.82
DeformTime 0.6705 0.5058 1.82 0.6664 0.4869 1.78 0.6638 0.4698 1.72 0.6669 0.4875 1.77
Sonnet 0.7098 0.4605 1.66 0.6890 0.4721 1.72 0.6856 0.4633 1.70 0.6948 0.4653 1.69
120 DLinear 0.6260 0.5535 1.99 0.4837 0.5678 2.07 0.6341 0.4773 1.75 0.5812 0.5328 1.94
iTransformer 0.6079 0.5563 2.00 0.6114 0.5013 1.83 0.5811 0.5124 1.88 0.6001 0.5233 1.90
PatchTST 0.6435 0.5017 1.81 0.4324 0.6084 2.22 0.6099 0.4972 1.82 0.5619 0.5357 1.95
TimeXer 0.6139 0.5558 2.00 0.6150 0.5044 1.84 0.5935 0.5042 1.85 0.6075 0.5215 1.90
Samformer 0.5653 0.5537 2.00 0.5719 0.5212 1.90 0.6136 0.4914 1.80 0.5836 0.5221 1.90
ModernTCN 0.6442 0.5410 1.95 0.5708 0.5321 1.94 0.5764 0.5233 1.92 0.5971 0.5321 1.94
Crossformer 0.6500 0.5573 2.00 0.6383 0.5133 1.87 0.6498 0.4929 1.81 0.6461 0.5212 1.89
DeformTime 0.6204 0.5384 1.94 0.6661 0.4815 1.75 0.6308 0.4951 1.81 0.6391 0.5050 1.83
Sonnet 0.6609 0.5005 1.80 0.6583 0.4794 1.75 0.6568 0.4692 1.72 0.6587 0.4830 1.76
D.9 Additional information for Sonnet’s ablation study
Information in this section supplements section 4.5 in the main paper.
$\neg$ Coher: This denotes that in MVCA, we do not compute the coherence attention. Instead of multiplying the attention with the value embedding ($\mathbf{O}_{h}=\mathbf{A}\odot\mathbf{V}_{h}$), the output embedding of each head $\mathbf{O}_{h}$ is equal to the value embedding, i.e. $\mathbf{O}_{h}=\mathbf{V}_{h}$. In this case, the attention highlighting temporal importance is not included; only the MLP that captures the variable dependency is included.
$\neg$ MLP: This denotes that we do not add the residual connection with the two-layer perceptron to obtain the attention output. Instead of computing $\mathbf{O}_{m}=\mathbf{O}_{r}+\texttt{MLP}\left(\mathbf{O}_{r}\right)$, we directly have $\mathbf{O}_{m}=\mathbf{O}_{r}$. In this case, the inter-variable dependency via the MLP is omitted. Although some inter-variable interactions are partially (no learnable parameter is included in this process) captured when transforming the input into the frequency space along the feature dimension.
$\neg$ MVCA: This denotes that we remove the MVCA from the modelling structure of Sonnet. After projecting the embedding into the wavelet space using the learnable wavelet transform, we directly feed the input into the Koopman evolvement layer.
$\neg$ Embed: This denotes that instead of embedding the endo- and exogenous variables separately with $2$ different weight matrices, we use $1$ single weight matrix, $\mathbf{W}\in\mathbb{R}^{(C+1)\times d}$, over all input variables, $\mathbf{Z}\in\mathbb{R}^{L\times(C+1)}$, to obtain the input embedding $\mathbf{E}\in\mathbb{R}^{L\times d}=\mathbf{Z}\mathbf{W}$.
$\neg$ Koop: This denotes that we do not predict the next step of the temporal state using the Koopman operator. Instead, after capturing the dependencies using MVCA, the time series is directly reconstructed from the output of the attention module.
Table S8: WEA forecasting results (temperature at $850$ hPa) in Cape Town (WEA-CT). We assess performance on $3$ test periods (years 2016, ’17, ’18) using linear correlation ($r$), MAE, and sMAPE ($\epsilon\%$). Best results are bolded and second best are underlined.
$H$ Model 2016 2017 2018 Average
$r$ MAE $\epsilon\%$ $r$ MAE $\epsilon\%$ $r$ MAE $\epsilon\%$ $r$ MAE $\epsilon\%$
4 DLinear 0.6899 3.0696 17.69 0.6677 3.3553 19.03 0.6757 3.0199 17.93 0.6778 3.1483 18.22
iTransformer 0.8466 2.1083 12.25 0.8325 2.3752 13.70 0.8464 2.0884 12.54 0.8418 2.1906 12.83
PatchTST 0.5931 3.3676 19.33 0.5788 3.7192 21.07 0.5372 3.4144 20.12 0.5697 3.5004 20.17
TimeXer 0.8464 2.1310 12.48 0.8412 2.4309 13.92 0.8407 2.1510 13.03 0.8428 2.2376 13.14
Samformer 0.8495 2.0911 12.19 0.8248 2.4305 14.03 0.8338 2.1580 13.04 0.8361 2.2265 13.09
ModernTCN 0.8955 1.7858 10.42 0.8852 1.9753 11.44 0.8806 1.8644 11.30 0.8871 1.8752 11.05
Crossformer 0.9114 1.5785 9.13 0.9070 1.7736 10.34 0.9135 1.5626 9.38 0.9106 1.6382 9.62
DeformTime 0.9044 1.7005 9.98 0.8880 1.9208 11.12 0.9057 1.6589 10.15 0.8994 1.7600 10.41
Sonnet 0.9169 1.5732 9.20 0.9068 1.7550 10.26 0.9160 1.5439 9.42 0.9133 1.6240 9.63
12 DLinear 0.3597 3.9731 22.60 0.3874 4.1974 23.32 0.3239 3.9091 22.89 0.3570 4.0265 22.94
iTransformer 0.4321 3.8589 22.11 0.4092 4.2382 23.80 0.3857 3.8253 22.58 0.4090 3.9741 22.83
PatchTST 0.3525 4.1364 23.67 0.3792 4.3989 24.52 0.3214 4.0377 23.78 0.3511 4.1910 23.99
TimeXer 0.4979 3.6755 20.94 0.4804 3.9628 22.14 0.4763 3.5340 20.86 0.4849 3.7241 21.31
Samformer 0.4245 3.9482 22.49 0.4137 4.2929 23.97 0.3626 3.8921 22.95 0.4003 4.0444 23.14
ModernTCN 0.4861 3.6737 21.11 0.4560 4.0876 22.78 0.4537 3.5671 21.10 0.4653 3.7761 21.67
Crossformer 0.5485 3.5536 20.27 0.5238 3.8275 21.42 0.5246 3.3985 20.16 0.5323 3.5932 20.62
DeformTime 0.5377 3.5610 20.29 0.5394 3.7613 21.00 0.5267 3.3821 20.10 0.5346 3.5681 20.46
Sonnet 0.5527 3.5423 20.27 0.5452 3.7261 20.89 0.5389 3.3613 19.99 0.5456 3.5432 20.38
28 DLinear 0.3869 3.9030 22.17 0.4155 4.1227 22.99 0.4047 3.7504 21.94 0.4024 3.9254 22.37
iTransformer 0.3651 4.0092 22.90 0.4111 4.1308 23.10 0.4143 3.7352 21.79 0.3969 3.9584 22.60
PatchTST 0.3824 3.9671 22.57 0.4173 4.1099 23.01 0.4097 3.7444 21.86 0.4031 3.9405 22.48
TimeXer 0.4448 3.8264 21.80 0.5026 3.9775 22.20 0.4052 3.6936 21.60 0.4509 3.8325 21.87
Samformer 0.3968 3.9169 22.27 0.4159 4.1299 23.09 0.4219 3.7249 21.76 0.4116 3.9239 22.37
ModernTCN 0.3972 3.9191 22.28 0.3855 4.2644 23.63 0.4480 3.6363 21.18 0.4102 3.9399 22.36
Crossformer 0.4491 3.8314 21.83 0.4861 3.9413 22.02 0.4135 3.6457 21.32 0.4495 3.8061 21.72
DeformTime 0.4704 3.7845 21.60 0.4990 3.9476 22.06 0.4510 3.5481 20.82 0.4735 3.7601 21.49
Sonnet 0.4829 3.7430 21.39 0.5017 3.8840 21.73 0.4436 3.5560 20.84 0.4761 3.7277 21.32
120 DLinear 0.2564 4.1508 23.51 0.4413 4.0823 22.74 0.2488 3.9379 22.90 0.3155 4.0570 23.05
iTransformer 0.3760 4.0470 22.88 0.4466 3.9902 22.30 0.3580 3.8046 22.30 0.3935 3.9473 22.49
PatchTST 0.3485 4.0798 23.27 0.4448 4.0213 22.37 0.2282 4.0629 23.42 0.3405 4.0547 23.02
TimeXer 0.4334 3.8432 21.84 0.4717 4.0044 22.36 0.4079 3.6759 21.44 0.4377 3.8412 21.88
Samformer 0.3890 4.0904 23.47 0.4684 3.9402 21.96 0.3858 3.6749 21.47 0.4144 3.9018 22.30
ModernTCN 0.4450 3.9583 22.94 0.4212 4.1221 23.01 0.2811 4.1861 24.68 0.3824 4.0889 23.54
Crossformer 0.4738 3.7648 21.45 0.5259 3.8475 21.54 0.3787 3.6852 21.49 0.4595 3.7659 21.49
DeformTime 0.4758 3.7615 21.44 0.4537 3.9905 22.23 0.3999 3.6602 21.34 0.4431 3.8040 21.67
Sonnet 0.4718 3.7594 21.41 0.5126 3.9088 21.86 0.4389 3.5438 20.79 0.4744 3.7373 21.35
Appendix E Supplementary results
This section supplements section 4 of the main paper.
E.1 Detailed forecasting results with the ILI and weather forecasting tasks
Complete results, broken down across all test periods, models, and forecasting horizons for the ILI rate prediction tasks in England (ILI-ENG), US Region 2 (ILI-US2), and US Region 9 (ILI-US9) are enumerated in Tables S1S2, and S3, respectively. Similarly, Tables S4S8 present the corresponding results for weather forecasting in London (WEA-LD), New York (WEA-NY), Singapore (WEA-SG), Hong Kong (WEA-HK), and Cape Town (WEA-CT). For both tasks, we report MAE, sMAPE ($\epsilon\%$), and the linear correlation coefficient (denoted by $r$) between the model predictions and the target values over each test period.
E.2 Supplementary experiment results with sMAPE score
Table S9: Performance (average sMAPE across the $4$ test seasons) of iTransformer, Samformer, and PatchTST on the ILI forecasting tasks (ILI-ENG/US2/US9) with different modifications to the naïve attention mechanism. ‘$\neg$ Attn’ denotes the removal of the residual attention module, and FNet / FED / VDAB refer to using the attention modules proposed in FNet [43], FEDformer [62], and DeformTime [43], respectively. This table supplements Table 2 that reports MAEs. Best results are bolded and second best are underlined.
Attention $H=7$ days $H=14$ days $H=21$ days $H=28$ days
ENG US2 US9 ENG US2 US9 ENG US2 US9 ENG US2 US9
iTransformer 26.38 23.24 18.57 36.67 28.17 22.44 48.93 30.03 24.11 55.35 36.75 31.05
$\neg$ Attn 26.22 25.52 21.22 37.10 30.54 25.77 54.52 34.77 29.89 61.97 41.18 34.97
FNet 30.16 26.99 22.27 41.28 31.12 26.29 57.16 36.48 29.69 64.26 41.06 35.00
FED 40.21 32.58 26.25 48.37 37.86 30.68 67.76 50.94 40.97 73.86 54.82 44.33
VDAB 27.46 24.55 19.21 38.10 30.39 23.29 48.19 32.37 23.07 54.16 35.39 26.97
MVCA 26.87 20.27 16.34 37.71 26.01 20.77 50.64 27.94 22.19 57.64 33.23 24.89
SamFormer 28.31 24.21 19.39 36.63 30.16 24.50 54.41 31.42 24.22 60.78 36.8 27.95
$\neg$ Attn 29.99 29.26 24.13 40.52 34.58 28.73 52.00 41.77 36.09 62.22 46.08 40.87
FNet 30.01 27.94 21.95 40.33 33.06 27.33 53.16 36.21 31.30 61.64 42.28 34.68
FED 40.18 32.35 26.34 48.91 37.65 30.84 67.79 49.99 41.26 73.17 53.91 44.85
VDAB 25.98 21.88 17.05 36.84 27.48 20.95 46.52 30.78 22.29 57.09 33.49 25.63
MVCA 26.13 20.85 16.90 38.14 26.14 21.06 49.15 28.55 22.57 55.27 32.39 24.66
PatchTST 27.61 24.51 19.34 37.76 30.11 24.09 51.11 36.70 29.40 59.60 42.61 33.35
$\neg$ Attn 28.54 26.99 21.51 41.64 31.82 26.36 53.64 38.75 31.75 60.11 43.25 36.17
FNet 31.93 28.81 22.14 44.15 33.77 25.64 58.17 38.11 26.46 60.51 41.48 29.92
FED 40.21 32.59 26.25 50.26 37.88 30.76 67.85 50.89 41.16 74.08 54.61 44.42
VDAB 26.88 22.99 17.99 38.08 30.99 20.64 50.49 32.45 23.90 60.16 34.89 25.61
MVCA 25.77 22.90 17.50 37.97 28.33 20.48 49.10 32.45 22.05 59.48 36.14 25.97
Results enumerated in Table S9 complement the ones presented in section 4.3, Table 2 by providing sMAPE estimates ($\epsilon\%$). On average, replacing the attention in the base models with MVCA reduces sMAPE by $9\%$. Moreover, compared to the best-performing variant (VDAB), our proposed MVCA reduces sMAPE by $2.3\%$. This reduction in relative error further consolidates our argument that MVCA provides better prediction accuracy in MTS forecasting.
E.3 Sensitivity to random seed initialisation
Table S10: Sensitivity to different seed initialisations ($5$ seeds) for Sonnet using the ILI-ENG data set across all forecasting horizons ($H$). $\mu$ and $\sigma$ denote the mean and standard deviation of the $5$ obtained MAEs per test period. The results in the main paper were obtained for seed ‘42’.
$H$ seed 2015/16 2016/17 2017/18 2018/19 Average
7 10 1.088 1.194 1.756 1.490 1.382
42 1.029 1.200 2.037 1.650 1.479
111 1.065 1.061 2.070 1.497 1.423
$1$ 1.188 1.182 1.815 1.220 1.351
1234 1.095 1.200 1.974 1.478 1.437
$\mu\leavevmode\nobreak\ (\sigma)$ 1.09 (0.06) 1.17 (0.06) 1.93 (0.14) 1.47 (0.16) 1.41 (0.05)
14 10 1.539 1.418 3.022 1.916 1.974
42 1.533 1.245 2.831 2.081 1.923
111 1.672 1.623 2.905 2.257 2.114
$1$ 1.590 1.492 2.937 2.049 2.017
1234 1.668 1.545 2.928 1.727 1.967
$\mu\leavevmode\nobreak\ (\sigma)$ 1.60 (0.07) 1.46 (0.14) 2.92 (0.07) 2.01 (0.20) 2.00 (0.07)
$H$ seed 2015/16 2016/17 2017/18 2018/19 Average
21 10 1.813 1.524 3.999 2.569 2.476
42 1.939 1.568 3.971 2.562 2.510
111 1.945 1.329 3.907 2.541 2.431
$1$ 1.966 1.487 4.001 2.469 2.481
1234 2.107 1.320 4.006 2.542 2.494
$\mu\leavevmode\nobreak\ (\sigma)$ 1.95 (0.10) 1.45 (0.11) 3.98 (0.04) 2.54 (0.04) 2.48 (0.03)
28 10 2.324 1.746 4.417 2.986 2.869
42 2.013 1.749 4.460 2.770 2.748
111 2.142 1.596 4.346 2.898 2.745
$1$ 2.322 1.599 4.606 2.955 2.870
1234 2.342 1.740 4.191 2.781 2.764
$\mu\leavevmode\nobreak\ (\sigma)$ 2.23 (0.15) 1.69 (0.08) 4.40 (0.15) 2.88 (0.10) 2.80 (0.06)
Results that assess the sensitivity of Sonnet to different random seed initialisations are enumerated in Table S10. We have examined $5$ different seeds (‘10’, ‘42’, ‘111’, ‘$1$’, ‘1234’) while conducting experiments on the ILI-ENG data set for all test periods and forecasting horizons. The obtained results highlight that the proposed model provides consistent results across all regions and forecasting horizons. Results do not deviate significantly from the ones presented in the main paper (Table 1) which are obtained using seed ‘42’. Standard deviations in the averaged MAEs (all test periods) over the $5$ seeds for each forecasting horizon do not exceed $0.07$ which in real terms (using the units of the actual data) represents $7$ people in a population of $10$ million. This highlights the robustness of the mean estimate.
E.4 Analysis of model performance for tasks with strong seasonality
Table S11: Effect of different $\alpha$ values for Sonnet while conducting experiments with the ELEC data set. $r$ and $\epsilon\%$ denote the correlation and sMAPE, respectively. “Naïve” in the last row denotes a seasonal persistence model. Best results are bolded and second best are underlined.
$\alpha$ 2020 2021
$H=12$ $H=24$ $H=36$ $H=12$ $H=24$ $H=36$
$r$ MAE $\epsilon\%$ $r$ MAE $\epsilon\%$ $r$ MAE $\epsilon\%$ $r$ MAE $\epsilon\%$ $r$ MAE $\epsilon\%$ $r$ MAE $\epsilon\%$
0.75 0.9793 0.1085 25.62 0.9748 0.1214 27.44 0.9614 0.1514 33.14 0.9780 0.1073 24.87 0.9733 0.1192 26.85 0.9634 0.1457 30.39
0.5 0.9807 0.1082 25.73 0.9747 0.1247 28.25 0.9671 0.1456 31.41 0.9782 0.1108 25.82 0.9747 0.1185 26.82 0.9640 0.1421 30.88
0.25 0.9813 0.1093 25.93 0.9749 0.1250 29.17 0.9680 0.1430 31.69 0.9788 0.1111 26.03 0.9735 0.1209 26.65 0.9654 0.1434 30.56
0 0.9823 0.1036 25.30 0.9772 0.1179 26.96 0.9694 0.1401 31.25 0.9806 0.1042 24.56 0.9753 0.1170 27.09 0.9679 0.1308 28.91
Naïve 0.9461 0.1675 34.30 0.9461 0.1675 34.30 0.9461 0.1675 34.30 0.9412 0.1736 35.14 0.9412 0.1736 35.14 0.9412 0.1736 35.14
Prior work has shown that modelling inter-variable dependencies in time series problems may occasionally cause overfitting [34]. While our results for the ILI and WEA forecasting tasks do not corroborate this (section 4.2), we conduct further experiments on a different task (ELEC) that displays a different pattern of performance across most models. In these experiments, we explore different values for the parameter $\alpha$ in Sonnet. $\alpha$ controls the relative contribution of exogenous variables in the learned representations (section 3.1). For $\alpha=0$, Sonnet only uses historical values of the target (endogenous) variable. We use the ELEC data set for evaluation because for this specific task, PatchTST—that does not capture inter-variable dependencies—outperforms some models that do (e.g. ModernTCN and Crossformer). In addition, $7$ out of $9$ evaluated models cannot surpass the accuracy of the seasonal persistence model (both in terms of MAE and sMAPE), indicating that for ELEC, seasonal patterns are consistent over time. Sonnet’s hyperparameters are re-tuned (see section 4.2 and  Appendix D), except for $\alpha$. Results for $\alpha=\{0,0.25,0.5,0.75\}$ are enumerated in Table S11. The last row contains the performance of the seasonal persistence model (denoted as “Naïve”) with a seasonality set to $168$ time steps ($1$ week).
We observe that although the prediction difference across different values of $\alpha$ is relatively small, Sonnet is consistently more accurate when setting $\alpha=0$, i.e. when only historical values of the target variable are used and no exogenous variables are included. This indicates that for ELEC the exogenous inputs are not very informative. The performance of the seasonal persistence model also supports this conclusion given its strong performance for both years 2020 and 2021 ($r>0.94$). This suggests that when strong seasonality is present, Sonnet benefits from following an autoregressive forecasting design without incorporating inter-variable dependencies (as $\alpha$ is a learnable parameter).
E.5 Computational cost of Sonnet
Figure S1: Computational cost (GPU memory) of Sonnet regarding the hidden dimension size $d$, the number of atoms $K$, the number of exogenous variables $C$, and the look back window length $L$.
We analyse Sonnet’s computational cost w.r.t. $4$ tunable hyperparameters: the number of atoms $K$, the size of the hidden dimension $d$, the length of the look-back window $H$, and the number of exogenous variables $C$. We analyse the influence of each variable independently by keeping the other three parameters fixed. Specifically, we set $d\in\{32,64,128,256\}$ while having $K=32$, $L=112$, $C=128$; similarly, we evaluate $K\in\{16,32,64,128\}$, $L\in\{56,84,112,224\}$, and $C\in\{64,128,256,512\}$, while the other parameters are fixed to $d=64$, $K=32$, $L=112$, and $C=128$, respectively. The computational cost of each setup is obtained by averaging memory consumption over $100$ epochs.
The GPU memory consumption of Sonnet during training is depicted in Figure S1. Sonnet’s GPU memory consumption stays constant as $C$ increases. This is because the input is mapped to a fixed embedding space for modelling. For the other three parameters, the method shows a linear increase in GPU memory consumption.
Appendix F Limitations
The main limitation is that the validity and predictiveness of the proposed method are only verified empirically (through experiments) but not theoretically. To mitigate against this, we have conducted extensive experiments over a wide range of data sets and tasks, covering different application fields. Many ($8$) of the data sets involve multiple test seasons (hence, multiple training / testing trials can be conducted), and have training sets that encompass time spans longer than a decade. Nevertheless, we acknowledge that additional evaluation across a broader range of forecasting problems could have provided an even more comprehensive assessment.
As far as the MVCA module is concerned (a key contribution of Sonnet), we argue that it is a direct and effective replacement for naïve attention in time series forecasting models. Although our current experiments support this claim, we nevertheless acknowledge that further analysis is needed to better understand its impact across a wider range of applications. Such insights could help guide more informed replacements of attention mechanisms in future work.
Finally, Sonnet—and many other SOTA forecasting models—cannot be considered as an interpretable model. We cannot know why Sonnet makes a certain forecast and we cannot directly attribute this to specific parts of the endogenous input. This can only be approximated through meta-analysis.
Appendix G Social Impact
Although it is difficult to predict the full scope of Sonnet’s influence, we can anticipate both beneficial and potentially adverse social implications arising from its adoption. On the positive side, enhanced time series forecasting—especially when grounded in realistic, application-oriented evaluation protocols—can substantially improve decision-making across a wide array of domains, including public health, energy demand and finance. By providing reliable projections, our approach may lead to more efficient use of scarce resources, better anticipation of system stresses and, ultimately, improved societal welfare. Furthermore, rigorous benchmarking against real-world scenarios helps to bridge the gap between academic research and practical deployment, encouraging translational collaborations that can accelerate the uptake of forecasting innovations in industry and government.
Conversely, over-reliance on algorithmic forecasts carries inherent risks. No matter how sophisticated, forecasting models (including Sonnet) cannot fully account for unforeseen systemic shocks, data biases or shifts in underlying processes. If taken as infallible oracles, such predictions may engender misplaced confidence, leading to suboptimal or even harmful policy and operational choices. To mitigate these dangers, it is essential that model outputs be interpreted judiciously by domain experts, who can contextualise quantitative predictions within broader qualitative insights.
Appendix H Supplementary ILI rate forecasting figures
In this section, a series of figures is shown below for the ILI rate prediction task, where each figure corresponds to a particular forecasting window and location, showcasing outputs from all models in parallel for comparison.
Figure S2: 7-day ahead forecasts for all influenza seasons and models, England (ILI-ENG).
Figure S3: 14-day ahead forecasts for all influenza seasons and models, England (ILI-ENG).
Figure S4: 21-day ahead forecasts for all influenza seasons and models, England (ILI-ENG).
Figure S5: 28-day ahead forecasts for all influenza seasons and models, England (ILI-ENG).
Figure S6: 7-day ahead forecasts for all influenza seasons and models, US Region 2 (ILI-US2).
Figure S7: 14-day ahead forecasts for all influenza seasons and models, US Region 2 (ILI-US2).
Figure S8: 21-day ahead forecasts for all influenza seasons and models, US Region 2 (ILI-US2).
Figure S9: 28-day ahead forecasts for all influenza seasons and models, US Region 2 (ILI-US2).
Figure S10: 7-day ahead forecasts for all influenza seasons and models, US Region 9 (ILI-US9).
Figure S11: 14-day ahead forecasts for all influenza seasons and models, US Region 9 (ILI-US9).
Figure S12: 21-day ahead forecasts for all influenza seasons and models, US Region 9 (ILI-US9).
Figure S13: 28-day ahead forecasts for all influenza seasons and models, US Region 9 (ILI-US9).